### v1

# Dual Conditional Diffusion Models for Sequential Recommendation

Abstract.
Recent advancements in diffusion models have shown promising results in sequential recommendation (SR). However, current diffusion-based methods still exhibit two key limitations. First, they implicitly model the diffusion process for target item embeddings rather than the discrete target item itself, leading to inconsistency in the recommendation process. Second, existing methods rely on either implicit or explicit conditional diffusion models, limiting their ability to fully capture the context of user behavior and leading to less robust target item embeddings. In this paper, we propose the Dual Conditional Diffusion Models for Sequential Recommendation (DCRec), introducing a discrete-to-continuous sequential recommendation diffusion framework. Our framework introduces a complete Markov chain to model the transition from the reversed target item representation to the discrete item index, bridging the discrete and continuous item spaces for diffusion models and ensuring consistency with the diffusion framework. Building on this framework, we present the Dual Conditional Diffusion Transformer (DCDT) that incorporates the implicit conditional and the explicit conditional for diffusion-based SR. Extensive experiments on public benchmark datasets demonstrate that DCRec  outperforms state-of-the-art methods.
Equal contribution.
copyright: nonecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Sequential recommender systems (SRSs) are designed to produce recommendations by predicting the next item that will capture a user’s interest (Wang et al., 2019a). Despite their effectiveness, existing models represent items with fixed vectors, which limit their ability to capture the latent aspects of items as well as the diverse and uncertain nature of user preferences (Li et al., 2024b; Yang et al., 2023b). Besides, these methods assume that the items with the most user interactions are the most relevant, which puts less-interacted items at a disadvantage and leads to the exposure problem (Yang et al., 2023b). Diffusion models (Ho et al., 2020; Ho and Salimans, 2022a), as an emergent deep generative model, add random noise to input data in a forward process and then recover the original data through a step-by-step reverse denoising process. Their denoising nature aligns well with the SR task, as modeling sequential item recommendations mirrors this step-wise approach (Xie et al., 2024). Recent studies on diffusion models for SR (Yang et al., 2023b) follow the paradigm of diffusion models, perturbing the embedding of the target item through a forward diffusion process into a Gaussian prior distribution. Then, they restore the Gaussian distribution iteratively through a reverse denoising process, also referred to as the sampling phase, to recover target representations and recommend items that are most similar to it.
Although these methods have achieved significant performance, current diffusion models for SR are still undergoing some shortcomings. Firstly, existing approaches (Yang et al., 2023b) commonly conduct conditional diffusion-based recommendations by conditioning the reverse denoising process on user behavior sequences. These methods focus on optimizing the diffusion process for the target item embeddings or generating an oracle item embedding for recommendation (Lin et al., 2024b). However, they do not explicitly model the diffusion process for the discrete target item itself (Gap 1). As illustrated in Figure 1(a), the traditional methods overlook a critical step: mapping the reversed target item representation into the discrete item index space. These methods often determine the recommended item by calculating the similarity (e.g., inner product) between the reversed target item representation and candidate item embeddings, selecting the item with the highest similarity score. While this method works for item ranking, it does not align with the core principles of diffusion models. Diffusion models are inherently probabilistic and continuous, but this step—based on a direct inner product comparison—introduces a deterministic, discrete decision process that breaks the continuous generative flow of diffusion. This inconsistency creates a gap between the optimization directions of the ranking loss for the recommendation task and the denoising loss for the diffusion model, resulting in suboptimal recommendation performance.
Figure 1. Figures (a) and (b) illustrate the differences between previous methods and our proposed approach in Markov modeling, while Figures (c) and (d) highlight the contrast between implicit and explicit conditional diffusion models for SR.
More importantly, existing diffusion-based SR methods can be broadly classified into two categories: implicit and explicit conditional approaches. As shown in Figure 1(c) and (d), in this paper, Implicit conditional methods refer to treating the user’s historical behavior as an implicit feature and integrating them into the generating target item at each diffusion step (Yang et al., 2023b; Li et al., 2024a, b). In contrast, Explicit conditional methods refer to leverage those historical behaviors as explicit guidance that directly influences the denoising process in diffusion SR (Lin et al., 2024a). The main difference between the two types of conditioning is that the implicit one can be regarded as a feature fusion with fixed and compressed historical features while the explicit one iteratively interacts with the generating target according to the diffusion progress. However, both approaches have inherent limitations. Implicit conditioning may oversimplify complex user behaviors by compressing them into a single vector, potentially losing valuable sequential information. Explicit conditioning, while providing detailed conditioning, can be more sensitive to noise and irrelevant actions within the user’s behavior sequence, which may hinder the recommendation quality (He and McAuley, 2016a). As no existing work combines implicit and explicit conditioning to leverage their respective strengths for enhanced SR (Gap 2). Another problem for existing diffusion models-based SR is their inefficient and slow sampling process, which relies on long Markov chains. Generating a target item requires iteratively refining random noise through numerous denoising steps, making the process computationally intensive and time-consuming (Lin et al., 2024b).
Aiming at bridging the aforementioned two significant gaps in existing SRSs, we propose the Dual Conditional Diffusion Models for SR (DCRec). Firstly, we begin with introducing the discrete-to-continuous sequential diffusion framework for SR. Our approach maintains consistency with the existing diffusion framework by introducing a Markov chain to model the transition from the reversed target item representation to the discrete item index. This probabilistic step fits more naturally within the diffusion process, allowing us to better capture the stochastic nature of item generation and selection. By explicitly modeling this transition, our method not only aligns with the underlying diffusion model but also provides a more coherent and accurate recommendation process. Furthermore, instead of following the existing implicit conditional or explicit conditional diffusion model-based SR methods, we propose novel conditioned diffusion models that integrate both implicit and explicit conditioning strategies. Based on our sequential recommendation diffusion framework, we introduce the Dual Conditional Diffusion Transformer (DCDT). This model concatenates historical information as an implicit condition while leveraging a cross-attention mechanism to incorporate explicit control signals throughout the diffusion process. By embedding dual conditions into both the forward and reverse diffusion stages, our approach ensures more precise and contextually relevant recommendations. In addition, the dual conditional mechanism leads to improved model performance, allowing the inference process to achieve optimal results with only a few sampling steps. This reduction in sampling steps significantly lowers computational overhead, making our model more efficient and better suited for real-time applications. The main contributions of this work are threefold:
  • We propose the Dual Conditional Diffusion Models for the sequential recommendation, introducing a novel diffusion framework for the SR. We explicitly introduce the discrete target items into the continuous diffusion process, defining the unified optimization target that considers both the ranking and the diffusion task.
  • Based on this framework, we propose a unified module, named Dual Conditional Diffusion Transformer (DCDT), which can generate accurate target item embedding based on both implicit and explicit conditions. Furthermore, our framework simplifies the diffusion process by enabling few-step inference instead of the traditional full-step inference.
  • We conduct extensive experiments on various public benchmark datasets and show that DCRec  outperforms the state-of-the-art SR models.
2. Related Work
$\bullet$ Diffusion Models. Diffusion models, derived from non-equilibrium thermodynamics, have achieved notable success in areas like computer vision, sequence modeling, and audio processing (Yang et al., 2024). Two primary paradigms exist unconditional models and conditional models. Unconditional models, such as DDPM (Ho et al., 2020), aim to push the performance boundaries of generative models. Conditional models incorporate guidance from labels to control generated samples. SGGM (Yang et al., 2022) conditions on self-produced labels, while You et al. (You et al., 2023) demonstrate how few labels and pseudo-training enhance both large-scale diffusion models and semi-supervised learners. Dhariwal (Dhariwal and Nichol, 2021) uses classifier guidance to improve sample quality, and Ho and Salimans (Ho and Salimans, 2022b) combine conditional and unconditional scores to balance sample quality and diversity. Multi-modal guidance further enriches the diffusion process, as seen in DiffuSeq (Gong et al., 2023) for NLP tasks and SDEdit (Meng et al., 2022) for image translation, with latent diffusion models (LDM) (Rombach et al., 2022) offering flexible, unified semantics. unCLIP (Ramesh et al., 2022) and ConPreDiff (Yang et al., 2023a) integrate CLIP latents in text-to-image generation.
$\bullet$ Diffusion Model-based Sequential Recommendation. Diffusion models show promise in the sequential recommendation by using user behavior sequences as guidance in the denoising process (Xie et al., 2024; Yang et al., 2023b; Li et al., 2024b). DiffuRec (Li et al., 2024b) corrupts target item embeddings with Gaussian noise and uses a Transformer to iteratively reconstruct them. DreamRec (Yang et al., 2023b) generates the oracle next-item embeddings based on user preferences but struggles with scalability due to its lack of negative sampling. DCDR (Lin et al., 2024a) adopts a discrete diffusion framework for progressive re-ranking. However, current methods face two main issues: they focus on embedding diffusion rather than discrete item generation, leading to inconsistencies, and they rely on implicit or explicit conditioning, limiting the capture of user context and weakening target item embeddings.
3. Preliminary
3.1. Problem Formulation
Let $\mathbb{U}=\{u_{1},u_{2},...,u_{|U|}\}$ and =$\mathbb{Z}=\{z_{1},z_{2},...,z_{|\mathbb{Z}|}\}$ be the user set and the item set. For each user $u\in\mathbb{U}$, there is a chronological sequence of historical interaction items denoted as $H_{z}$, where $H_{z}\subsetneqq\mathbb{Z}$. Generally, for a user $u$, given their sequence context $H_{z}=\{z_{1},z_{2},...,z_{n-1}\}$, the task of an SRS is to predict the next item (target item) $z$ which may interest $u$.
3.2. Diffusion Models for Continuous Domains
Recently, diffusion models have already demonstrated significant performance in continuous domains like computer vision and audio (Ho et al., 2020; Ho and Salimans, 2022a; Kong et al., 2021; Rombach et al., 2022). Thus, before presenting our diffusion SR model, we briefly recap the diffusion models in continuous domains (Ho et al., 2020; Luo, 2022).
A generative model is expected to match estimation distribution $p_{\theta}(e_{0})$ to the ground truth distribution $p(e_{0})$ as closely as possible, where $\theta$ represents learnable parameters and $e_{0}\in\mathbb{R}^{d}$ is a set of samples. For vanilla continuous diffusion models, they model the sample $e_{0}$ as a Markov chain with a sequence of latent variables $\{e_{T},\dots,e_{t},e_{t-1},\dots,e_{0}\}$, where $T$ is the diffusion step and $e_{T}$ is a standard Gaussian, $e_{t}$ is the intermediate latent variable. By this means, the latent dimension is exactly equal to the data dimension of $e_{0}$, and the diffusion model progressively recovers $e_{0}$ from Gaussian noise. Each denoising transition $e_{t}\rightarrow e_{t-1}$ is parameterized by a linear Gaussian transition, calculated by a denoising DNN $f_{\theta}(e_{t},t)$:
(1) $p_{\theta}(e_{t-1}|e_{t})=\mathcal{N}(e_{t-1};\mu_{\theta}(e_{t},t),\Sigma_{\theta}(e_{t},t))$
Thus, the joint distribution of diffusion models can be written as:
(2) $\small p_{\theta}(e_{0:T})=p_{\theta}(e_{T})\prod^{T}_{t=1}p_{\theta}(e_{t-1}|e_{t})$
(3) $where,p_{\theta}(e_{T})=\mathcal{N}(e_{T};\mathbf{0},\mathbf{I})$
To train the denoising model $f_{\theta}(e_{t},t)$, diffusion models first construct a progressive noising process $\{e_{0},\dots,e_{T-1},e_{T}\}$, known as the forward process. Corresponding to this, the denoising process is called the reverse process. The forward process adds Gaussian noise to $e_{0}$ step-by-step until the final latent $e_{T}$ is guaranteed to a standard Gaussian. Each noising transition $e_{t-1}\rightarrow e_{t}$ is defined as a linear Gaussian transition $q(e_{t}|e_{t-1})=\mathcal{N}(e_{t};\sqrt{\alpha_{t}}e_{t-1},(1-\alpha_{t})\mathbf{I})$, where the Gaussian parameters $\alpha_{t}\in(0,1)$ varies over time step $t$ and controls the noise scale added to the embedding in the current step $t$. Due to the Markov property and the Bayes’ theorem, the transition can be rewritten with $e_{0}$ as:
(4) $q(e_{t}|e_{t-1})=q(e_{t}|e_{t-1},e_{0})=\frac{q(e_{t-1}|e_{t},e_{0})q(e_{t}|e_{0})}{q(e_{t-1}|e_{0})}$
To optimize the parameter $\theta$, the diffusion model is trained to maximize the log-likelihood of $p_{\theta}(x)$ according to observed samples $e_{0}$. This objective can be formulated as $\arg\max_{\theta}\log p_{\theta}(e_{0})$. Mathematically, we can derive a proxy objective called the Evidence Lower Bound (ELBO), a lower bound of the likelihood term $\log p_{\theta}(e_{0})$. Formally, the equation of the ELBO is:
(5) $\mathbb{E}_{q(e_{1:T}|e_{0})}\left[\log\frac{p_{\theta}(e_{0:T})}{q(e_{1:T}|e_{0})}\right]$
Rather than directly maximize the likelihood term, diffusion models minimize the minus ELBO as (Luo, 2022; Ho et al., 2020):
(6) $\arg\max\limits_{\theta}\log p_{\theta}(e_{0})$ $\approx\arg\max\limits_{\theta}\mathbb{E}_{q(e_{1:T}|e_{0})}\left[\log\frac{p_{\theta}(e_{0:T})}{q(e_{1:T}|e_{0})}\right]$
(7) $=\arg\min\limits_{\theta}\mathbb{E}_{q(e_{1:T}|e_{0})}\left[\log\frac{q(e_{1:T}|e_{0})}{p_{\theta}(e_{0:T})}\right]$
Combining with Equations 2 and 4, the log term in Equation 7 can be further derived as:
(8) $\log\frac{q(e_{T}|e_{0})}{p_{\theta}(e_{T})}+\sum_{t=2}^{T}\log\frac{q(e_{t-1}|e_{t},e_{0})}{p_{\theta}(e_{t-1}|e_{t}))}-\log p_{\theta}(e_{0}|e_{1})$
Therefore, based on Equations 7 and 8, the ELBO-based optimization object of parameter $\theta$ is to minimize:
(9) $\mathcal{L}_{\text{elbo}}(e_{0})=\mathbb{E}_{q(e_{1:T}|e_{0})}$ $\Big{[}\log\frac{q(e_{T}|e_{0})}{p_{\theta}(e_{T})}+\sum_{t=2}^{T}\log\frac{q(e_{t-1}|e_{t},e_{0})}{p_{\theta}(e_{t-1}|e_{t})}$
$-\log p_{\theta}(e_{0}|e_{1})\Big{]}$
To simplify and combine the second and third terms in Equation 9, recent research (Luo, 2022) derivative a simple surrogate objective to obtain a mean-squared error term:
(10) $\mathcal{L}_{\text{elbo}}(e_{0})=\mathbb{E}_{q(e_{1:T}|e_{0})}\left[\sum_{t=1}^{T}\|f_{\theta}(e_{t},t)-e_{0}\|^{2}\right]$
4. Dual Conditional Diffusion Models for Sequential Recommendation
Figure 2. (a) illustrates a simplified workflow of our DCRec. During the end-to-end diffusion process, DCRec  implicitly concatenates the noised target item embedding $\bm{e_{t}}$ and the corresponding noised history embedding sequence $\bm{H_{t}}$ at each diffusion step. Meanwhile, DCRec  is also guided by clear history embedding $\bm{H_{0}}$ as explicit conditional signals. (b) depicts the forward and reverse diffusion process with conditioning. There are two denoising trajectories with implicit guidance. The dual conditional trajectory with additional explicit guidance is more likely to reach the target item (the orange area).
In this section, we present the comprehensive introduction of our DCRec  as shown in Figure 2, which includes both the Sequential Recommendation Diffusion Framework and the Dual Conditional Diffusion Transformer (DCDT). Firstly, we expend the vanilla continuous diffusion to the discrete sequential recommendation task and provide a theoretical-grounded diffusion framework. Then, we introduce the DCDT, our denoising network $f_{\theta}$ for diffusion SR under the proposed framework. Finally, we focus on the efficient inference for the DCRec.
4.1. Sequential Recommendation Diffusion Framework
Continuous diffusion models have achieved remarkable success in domains like vision and audio (Kong et al., 2021), but their application to SR has been limited due to the inherently discrete nature of the recommended items. Existing diffusion models in SR typically determine the recommended item by computing the similarity (e.g., inner product) between the reversed target item representation and candidate item embeddings, selecting the item with the highest similarity score. While this approach is effective for ranking, it fundamentally conflicts with the principles of diffusion models. Vanilla diffusion models are designed to be probabilistic and continuous. Current diffusion SR approaches (Yang et al., 2023b; Li et al., 2024b; Xie et al., 2024) directly apply the continuous diffusion framework to the SR task neglecting the discrete-to-continuous transition. This breaks the transition Markov chain, the generative flow that diffusion models rely on, creating a disconnect between the ranking objective and the diffusion model’s denoising process. Specifically, this misalignment leads to an inconsistency between the optimization directions of the ranking loss for recommendation and the denoising loss of the diffusion process, resulting in sub-optimal recommendation performance.
To address this issue and motivated by diffusion models in text domains (Li et al., 2022), we explicitly extend continuous diffusion models to discrete item domains as shown in Figure 2(a). As for discrete target item $z$ from the item pool $\mathbb{Z}$, the Markov chain in the forward and reverse processes are extending as $\{z,\dots,e_{T-1},e_{T}\}$ and $\{e_{T},\dots,e_{0},z\}$. Specifically, to map the discrete variables into continuous domains, we define a learnable embedding function $\text{Emb}(z)$. The forward transition process of $z$ is defined as:
(11) $q_{\phi}(e_{0}|z)=\mathcal{N}(e_{0};\text{Emb}(z),\mathbf{0}),$
where $\phi$ represents the learnable parameters in $\text{Emb}(z)$. As for the reverse process, we define the predicted distribution of $z$ as:
(12) $p_{\phi}(z|e_{0})=\arg\max\limits_{z}\text{cos}(e_{0},\text{Emb}(\mathbb{Z})),$
where $\text{cos}(\cdot)$ is the cosine similarity between $e_{0}$ and each item embedding of $\text{Emb}(\mathbb{Z})$. However, the transition distribution $p_{\phi}(z|e_{0})$ lacks a direct analytical formula.
After defining $p_{\phi}$ and $q_{\phi}$, we extend the ELBO in Equation 7 to include discrete item $z$ as:
(13) $\arg\max\limits_{\theta}\log p_{\theta}(z)$ $\approx\arg\min\limits_{\phi,\theta}\mathbb{E}_{q_{\phi}(e_{0:T}|z)}\left[\log\frac{q_{\phi}(e_{0:T}|z)}{p_{\phi,\theta}(z,e_{0:T})}\right],$
and the training objective in Equation 9 is rewritten as:
(14) $\mathcal{L}_{\text{elbo}}(z)=\mathbb{E}_{q_{\phi}(e_{0:T}|z)}\left(\mathcal{L}_{T}+\mathcal{L}_{t}+\mathcal{L}_{z}\right),\rm where\$
[…]
There are three terms of objectives in $\mathcal{L}_{\text{elbo}}$: the prior matching term $\mathcal{L}_{T}$, the denoising term $\mathcal{L}_{t}$ and the domain transition term $\mathcal{L}_{z}$. The prior matching term $\mathcal{L}_{T}$ is similar to the same term in Equation 9, which requires the forward process to transit the $e_{0}$ into Gaussian noise $p_{\theta}(e_{T})$. Continuous diffusion models (Luo, 2022; Ho et al., 2020; Ho and Salimans, 2022a; Kong et al., 2021) omit this term since there are no associated trainable parameters. This term is essential in $\mathcal{L}_{\text{elbo}}$ due to the correlation among $z$, $e_{0}$ and $\phi$. This term is equivalent to minimizing $\|e_{0}\|^{2}$. As the second term $\mathcal{L}_{t}$ require $f_{\theta}(e_{t},t)$ and $e_{0}$ as close as possible, we transfer the objective $\|e_{0}\|^{2}$ to $\|\sum_{t=1}^{T}f_{\theta}(e_{t},t)\|^{2}$ for fully optimize the DNN models.
The embedding entropy term $\mathcal{L}_{z}$ can be regarded as a KL-divergence $D_{\text{KL}}(q_{\phi}(e_{0}|z)||p_{\phi}(z|e_{0}))$. Although $q_{\phi}(e_{0}|z)$ is a linear transition, the distribution of $p_{\phi}(z|e_{0})$ is unknown. Therefore, we replace $D_{\text{KL}}(q_{\phi}(e_{0}|z)||p_{\phi}(z|e_{0}))$ by $D_{\text{KL}}(q(z)||p_{\phi}(z))=D_{\text{KL}}(q(z)||p_{\phi}(z))$. The simplified objective of $\mathcal{L}_{\text{elbo}}$ is:
(16) $\mathcal{L}_{\text{elbo}}(z)=\mathbb{E}_{q_{\phi}(e_{0:T}|z)}\left(\mathcal{L}_{T}+\mathcal{L}_{t}+\mathcal{L}_{z}\right),\rm where\$
(17) $\mathcal{L}_{T}$ $=\|f_{\theta}(e_{t},t)\|^{2},$
(18) $\mathcal{L}_{t}$ $=\|f_{\theta}(e_{t},t)-e_{0}\|^{2},$
(19) $\mathcal{L}_{z}$ $=D_{\text{KL}}(q_{\phi}(z)||p_{\phi}(z))$
4.2. Dual Conditional Diffusion Transformer
In this section, we provide a detailed explanation of the Dual Conditional Diffusion Transformer (DCDT) in DCRec, which is a robust denoising transformer built within the framework in Section 4.1. Previous methods in diffusion-based SR have primarily employed either implicit or explicit conditioning mechanisms to incorporate user behavioral data into the recommendation process (Yang et al., 2023b; Li et al., 2024a; Xie et al., 2024). Almost implicit conditional approaches compress the user’s historical behavior sequence into a compact representation vector, which is then used to guide the reverse diffusion process (Lin et al., 2024b; Li et al., 2024a). While this method efficiently captures global user preferences, it may inadvertently oversimplify complex user behaviors, losing important sequential and temporal dynamics inherent in the user’s interactions. This compression can lead to a loss of valuable information about the order and context of past actions, which are crucial for accurate SRs. Explicit conditioning methods (Lin et al., 2024a), on the other hand, skip the feature compression step and directly introduce the entire sequence of historical item embeddings during the reverse generation process. This allows the model to access detailed step-by-step interactions and capture fine-grained sequential dependencies. However, explicit conditioning on the whole historical information, these methods can be more sensitive to noise and irrelevant actions within the user’s behavior sequence, which may hinder the recommendation quality (He and McAuley, 2016a, b).
To tackle these limitations, we propose the Dual Conditioned Diffusion Transformer (DCDT) as shown in Figure 3, a novel approach that integrates both implicit and explicit conditioning strategies. By combining the representation of the user’s global preferences with the direct utilization of detailed historical behaviors, DCDT  leverages the strengths of both methods. The implicit component captures global user interests and long-term behavioral patterns, providing a stable guide for the diffusion process. Simultaneously, the explicit component preserves sequential complexities and temporal dynamics by referencing specific past interactions during generation. DCDT  employs a dual conditional diffusion process that incorporates an implicit conditional historical forward process and an explicit conditional reverse process, leading to more accurate recommendations. As shown in Figure 2(b), the dual conditional trajectory (green line) in the reverse diffusion process merges these two guiding signals, preventing the denoising trajectory from being overly influenced by noise or irrelevant historical items. This effectively narrows the trajectory toward the correct target item, improving the precision of the recommendation.
4.2.1. Implicit Conditional Historical Forward Process:
Previous implicit conditional methods tend to encode the user’s historical behavior sequences into compact feature representations, which are then used to guide the reverse diffusion process (Yang et al., 2023b; Li et al., 2024a). However, compressing the historical behavior sequences into a single vector might lead to the loss of important sequential information due to the sparse nature of SRSs (He and McAuley, 2016b). In our approach, we avoid this compression step. Instead, we combine the uncompressed historical behavior sequences with the target item and add noise to this combined input before feeding it into the diffusion model. We believe that this allows the diffusion model to generate a more robust representation of the target item, as it can better capture detailed sequential patterns and contextual information from the user’s history. For convenience, for our implicit conditional diffusion models, we use $S_{0}$ to donate the concatenated sequence of the history item sequence $H_{0}$ and the target $e_{0}$, where $S_{0}=\text{Concat}(H_{0},e_{0})$. Thus, we continuously incorporate Gaussian noises into the whole item sequences $S_{0}$ rather than the target item embedding $e_{0}$ as follows:
(20) $q(S_{t}|S_{t}-1)$ $=\mathcal{N}(S_{t};\sqrt{\alpha_{t}}S_{t-1},(1-\alpha_{t})\mathbf{I})$
(21) $=\sqrt{\alpha_{t}}S_{0}+\sqrt{1-\alpha_{t}}S_{t-1}$
Based on the superposition of Gaussian distribution and Equation 21, we can calculate the transition from $S_{0}$ to $S_{t}$ as:
(22) $q(S_{t}|S_{0})=\sqrt{\bar{\alpha}_{t}}S_{0}+\sqrt{1-\bar{\alpha}_{t}}S_{T},\ S_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, and $S_{t}$ denotes the item embeddings in the forward step $t$. We further add explicit conditioning by adding $H_{0}$ with a scaling factor $\delta$ to $S_{T}$ and rewrite Equation 22 as:
(23) $q(S_{t}|S_{0})=\sqrt{\bar{\alpha}_{t}}S_{0}+\sqrt{1-\bar{\alpha}_{t}}(S_{T}+\delta H_{0}),$
Therefore, in the ELBO-based training objective, we can rewrite Equation 17 and Equation 18 as:
[…]
Compared to existing implicit diffusion models-based SR methods (Yang et al., 2023b), our method adds noising the entire user sequence rather than just the target item. This approach allows the following reverse process to explicitly leverage the user’s interaction history when generating the target item that aligns with the user’s interests, leading to more accurate and relevant recommendations.
Figure 3. Our DCDT  comprises explicit conditional layer normalization, multi-head self-attention, and an explicit cross-attention.
4.2.2. Explicit Conditional Reverse Process:
In the context of SR, the restoration process reverses the transition modeling by gradually eliminating the noise and recovering the original sequential patterns of user behavior, thereby restoring the data distribution underlying the user’s item preferences. Meanwhile, cross attention as an explicit conditional method has proved effective in classic conditional generation tasks (Rombach et al., 2022; Podell et al., 2024). Therefore, we further introduce the user’s item preferences $H_{0}$ as an explicit conditional signal to the denoising transformer $f_{\theta}$ in Equation 24 as:
(25) $\hat{S}_{0}=f_{\theta}(S_{t},t,H_{0})$
Firstly, we introduce the Conditional Layer Normalization, abbreviated as CondLN. In our SR setting, we hope to extend traditional layer normalization to our explicit conditional denoising process. Therefore, we replace the element-wise learnable parameters gain $g$ and bias $b$ of affine transformation in the original Layer Normalization (Lei Ba et al., 2016) by integrating user interaction sequence $S_{0}$ and diffusion step embedding $t$. The CondLN can be formulated as:
(26) $\bar{S}_{t}=g^{\prime}(t,H_{0})\times\frac{S_{t}-\mathbb{E}(S_{t})}{\sqrt{\text{Var}(S_{t})+\epsilon}}+b^{\prime}(t,H_{0})$
where $\mathbb{E}(\cdot)$ and $\text{Var}(\cdot)$ denote the mean and standard variance of $S_{t}$, and $\epsilon$ denotes the small constant for numerical stability. $g^{\prime}(\cdot)$, $b^{\prime}(\cdot)$ are the MLP function conditioned by $H_{0}$ and $t$. We scale the normalized $\text{Var}(S_{t})$ and shift it with $\mathbb{E}(S_{t})$.
In this way, CondLN explicitly affects the generation process with user interaction history $H_{0}$. This allows for more adaptive and context-aware normalization compared to standard layer normalization, which is crucial for capturing the intricate dependencies in SR tasks. Next, we input this sequence embedding $\bar{S}_{t}$ into a self-attention module (SA):
(27) $\mathbf{\bar{S}_{t}}=\text{SA}(\bar{S}_{t})=\text{Softmax}\left(\frac{(\bar{S}_{t}\mathbf{W}^{Q})(\bar{S}_{t}\mathbf{W}^{K})^{\top}}{\sqrt{d}}\right)(\bar{S}_{t}\mathbf{W}^{V})\in\mathbb{R}^{(n-1)\times d}$
where $\mathbf{W}^{Q}$,$\mathbf{W}^{K}$,$\mathbf{W}^{V}$ are learnable weight parameters, $d$ denotes the dimension of embedding. After obtaining the self-attention output, then we leverage the Cross-attention layer (CA):
(28) $\bar{S}_{t}=\text{CA}(\bar{S}_{t},\mathbf{H_{0}})=\text{Softmax}\left(\frac{(\bar{S}_{t}\mathbf{W}^{Q})(\mathbf{H_{0}}\mathbf{W}^{K})^{\top}}{\sqrt{d}}\right)(\mathbf{H_{0}}\mathbf{W}^{V})$
where $\mathbf{W}^{Q}$, $\mathbf{W}^{K}$, $\mathbf{W}^{V}$ are learnable parameters. Due to the sparse interaction in SR (He and McAuley, 2016b), we adapt $\bar{S}_{t}$ as query and the $H_{0}$ (pure historical embeddings) as key, value. $H_{0}$ can provide more rich user interaction information than ambiguous $H_{t}$. Hence, we select a cross-attention architecture as the denoising decoder instantiation capable of taking the entire sequence representation and corresponding step indicator $t$ as input. Hence, we adopt a cross-attention architecture for the denoising decoder, which effectively processes the entire sequence representation along with the corresponding diffusion step indicator $t$ as input. Then we endow the traditional layer normalization and feed-forward network to obtain the final user sequence embedding $\hat{S}_{0}$:
(29) $\bar{S}_{t}=\text{LayerNorm}(\bar{S}_{t}+\text{Dropout}(\text{sublayer}(\bar{S}_{t})))$
(30) $\hat{S}_{0}=\text{FFN}(\bar{S}_{t})=\text{RELU}(\bar{S}_{t}\times\mathbf{W}+\mathbf{b})$
After obtaining the $\hat{S}_{0}$, we extract the last item embedding $e_{0}$ of $\hat{S}_{0}$ as the generative target item embedding. Then we leverage the variant of Equation 12, to transit continuous $\hat{S}_{0}$ into the discrete item space $\mathbb{Z}$ as $z$, taking the item index $z$ corresponding to the maximal as the final recommendation list to the specific user. Finally, the optimization object of the dual conditional diffusion process can be reformulated as:
(31) $\mathcal{L}_{\text{elbo}}(z)=\mathbb{E}_{q_{\phi}(e_{0:T}|z)}\left[\mathcal{L}_{T}+(1-\lambda)\mathcal{L}_{t}+\lambda\mathcal{L}_{z}\right],\rm where\$
(32) $\mathcal{L}_{T}$ $=\|f_{\theta}(e_{t},t)\|^{2},$
(33) $\mathcal{L}_{t}$ $=\|f_{\theta}(e_{t},t)-e_{0}\|^{2},$
(34) $\mathcal{L}_{z}$ $=D_{\text{KL}}(q_{\phi}(z)||p_{\phi}(z))$
where $\lambda$ donates the balance factor mediating the focus between $\mathcal{L}_{t}$ and $\mathcal{L}_{z}$. Algorithm 1 in Appendix A provides a detailed explanation of DCRec’s training process.
4.3. Model Inference and Inference Acceleration
In the Section 4.1 and Section 4.2, we have introduced the optimization process in DCRec  and the training of DCDT. The denoising model DCDT  is trained to provide predicted $\hat{S}_{0}$ to approximate $S_{0}$. Similar to Equation 2, the inference Markov transition of DCRec  can be formulated as:
(35) $p_{\theta}(S_{t-1}|S_{t})=\mathcal{N}(S_{t-1};\mu_{\theta}(\hat{S}_{0}(S_{t}),t),\Sigma_{\theta}(\hat{S}_{0}(S_{t}),t))$
This can be described as a repeated approximation-and-adjust process since each step aims to predict $\hat{S}_{0}(S_{t})$ and provide $S_{t-1}$, which is less noise than $S_{t}$, for the next step. The time cost of inference process linearly increases associated with the increase of the total number of steps. A natural idea for acceleration is that, if the approximation at the early is accurate enough with no need for adjustment, DCRec  can directly skip several intermediate steps. Here, we define the number of skipping steps as k. We rewrite Equation 35 as:
(36) $p_{\theta}(S_{t-k}|S_{t})=\mathcal{N}(S_{t-k};\mu_{\theta}(\hat{S}_{0}(S_{t}),t),\Sigma_{\theta}(\hat{S}_{0}(S_{t}),t))$
For our DCRec, we observe that DCDT  is robust enough to fulfill Equation 36 at an early stage, which significantly reduces time overhead. We will discuss the details in Section 5.4. For the pseudo-code of the inference, please refer to Algorithm. 2 in the Appendix.
5. Experiment
In this section, we conduct a comprehensive experimental study to address the following research questions (RQs):
$\bullet$ RQ1: How does the performance of DCRec  compare with other baselines across the datasets in different experiment settings?
$\bullet$ RQ2: What is the impact of different components (e.g., reconstruction loss, ranking loss, regularization loss, and DCDT  modules) within the DCRec  on overall performance?
$\bullet$ RQ3: How does DCRec  perform in terms of inference efficiency?
5.1. Experimental Settings
$\bullet$ Datasets. We evaluate our proposed DCRec  on three publicly available datasets: Amazon 5-core Beauty, Amazon 5-core Toys, and Yelp, following the experimental setup of previous works (Zhou et al., 2020a; Chen et al., 2022a; Li et al., 2024b). Detailed descriptions of these datasets can be found in Appendix B due to space constraints.
$\bullet$ Baselines. To evaluate the performance of our model, we select various baselines covering three types of mainstream methods in SR tasks, e.g., representative SR methods, contrastive learning SR methods, and diffusion-based SR methods. The detailed information is listed in the Appendix C.
$\bullet$ Implementation Details and Parameter Settings. To ensure fair comparisons, all the baselines and ours are implemented based on the popular recommendation framework RecBole (Zhao et al., 2021) under the same settings and evaluated consistently. The detailed information is listed in the Appendix F.
$\bullet$ Evaluation Metrics. We use two commonly used metrics including Normalized Discounted Cumulative Gain @K (NDCG@K), Hit Ratio @K (HR@K) to evaluate the performance of all compared methods. For all the baseline models, we generate the ranking list of items for each testing interaction. Both of them are applied with $K$ chosen from {5, 10}, evaluating the ranking results using all items in the dataset as candidates. Meanwhile, we follow the setting in (Wang et al., 2019b), using the paired t-test with p¡0.05 for the significance test.
5.2. Overall Performance (RQ1)
Table 1. Overall performance. Bold scores represent the highest results of all methods. Underlined scores stand for the second-highest results. Our model achieves the state-of-the-art result among all baseline models. means the improvement is significant at $p<0.05$, with the relative improvements denoted as Improv.
Dataset Metric GRU4Rec Caser SARSRec BERT4Rec S3Rec CL4SRec ICLRec DreamRec DiffuRec SdifRec DCRec Improv.
Beauty HR@5 0.0363 0.0374 0.0337 0.0209 0.0189 0.0423 0.0493 0.0498 0.0559 0.0609 $\textbf{0.0630}^{*}$ 3.45%
HR@10 0.0571 0.0601 0.0635 0.0393 0.0307 0.0694 0.0737 0.0698 0.0783 0.0819 $\textbf{0.0859}^{*}$ 4.88%
NDCG@5 0.0243 0.0241 0.0237 0.0110 0.0115 0.0281 0.0324 0.0325 0.0406 0.0437 $\textbf{0.0449}^{*}$ 2.74%
NDCG@10 0.0310 0.0332 0.0322 0.0169 0.0153 0.0373 0.0393 0.0397 0.0478 0.0507 $\textbf{0.0523}^{*}$ 3.16%
Toys HR@5 0.0286 0.0208 0.0432 0.0274 0.0143 0.0526 0.0590 0.0510 0.0550 0.0588 $\textbf{0.0690}^{*}$ 17.35%
HR@10 0.0440 0.0322 0.0660 0.0450 0.0094 0.0776 0.0830 0.0635 0.0739 0.0758 $\textbf{0.0903}^{*}$ 19.13%
NDCG@5 0.0179 0.0135 0.0293 0.0174 0.0123 0.0362 0.0403 0.0316 0.0411 0.0447 $\textbf{0.0518}^{*}$ 15.88%
NDCG@10 0.0228 0.0172 0.0347 0.0231 0.0391 0.0428 0.0479 0.0391 0.0471 0.0498 $\textbf{0.0587}^{*}$ 17.87%
Yelp HR@5 0.0142 0.0160 0.0162 0.0196 0.0101 0.0229 0.0257 0.0174 0.0364 0.0233 $\textbf{0.0405}^{*}$ 11.26%
HR@10 0.0248 0.0254 0.0311 0.0339 0.0176 0.0392 0.0426 0.0193 0.0551 0.0375 $\textbf{0.0611}^{*}$ 10.89%
NDCG@5 0.0080 0.0101 0.0096 0.0121 0.0068 0.0144 0.0162 0.0117 0.0253 0.0157 $\textbf{0.0272}^{*}$ 7.51%
NDCG@10 0.0124 0.0113 0.0136 0.0167 0.0092 0.0197 0.0217 0.0152 0.0313 0.0201 $\textbf{0.0338}^{*}$ 7.99%
In this section, Table 1 reports the comparison results between our method and 10 different baseline methods on three datasets. We can draw the following observations: (1) Our DCRec  achieves significant improvements across all datasets, confirming the effectiveness of our diffusion approach in modeling SR tasks. By introducing a complete Markov chain to model the transitions from the reversed target item representations to discrete item indices, our framework bridges the gap between discrete and continuous item spaces. This leads to improved consistency with the diffusion process, which is reflected in the performance gains over baseline methods. For example, our DCRec  outperforms SdifRec on the Beauty dataset with a 3.45% improvement in HR@5 and a 3.16% improvement in NDCG@10, demonstrating the robustness of our approach. (2) Generative models, including DiffuRec, DreamRec, and SdifRec, generally perform better than traditional SR and CL-based SR methods such as SARSRec and CL4SRec. This trend highlights the advantage of generative modeling techniques in capturing complex user-item relationships. Among these, diffusion-based methods, such as DiffuRec and SdifRec, tend to exhibit superior performance, likely because diffusion models can capture the uncertainty of user interest. However, DCRec  surpasses both DiffuRec and SdifRec on all datasets, demonstrating the added benefit of our dual conditional diffusion mechanism, which effectively captures both implicit and explicit conditions in the forward and reverse processes.
5.3. In-depth Analysis (RQ2)
5.3.1. Ablation Study.
We execute the ablation study for DCRec from two distinct angles: loss perspective and module perspective.
Figure 4. The HR@5 results of training DCRec  using different types of optimization losses on Beauty and Toys datasets.
Loss function perspective. We evaluate the distinct contributions of different loss components $L_{T}$ (Regularization loss), $L_{t}$ (Diffusion MSE loss), $L_{z}$ (Ranking loss) in Amazon Beauty and Toys datasets, which is depicted as Figure 4. Based on these results, we can find that (1) When optimizing with only $L_{T}$ or $L_{t}$, the model shows relatively lower performance compared to using $L_{z}$ on both datasets, indicating the ranking loss plays a more significant role in the SR task. (2) The combination of losses leads to significantly better performance. This demonstrates that combining regularization, diffusion MSE, and ranking losses results in a more robust optimization process, leading to better overall model performance. (3) Overall, while $L_{T}$, $L_{t}$ and $L_{z}$ losses jointly optimize our DCRec, our model can achieve the best performance.
Module perspective. To observe the impact of different components of DCDT , we conduct additional experiments on Amazon Beauty, Toys, and Yelp datasets. The results are shown in Figure 5.
Figure 5. The HR@5 results of using different modules.
The results compare four configurations: single attention (Single-attn) with and without Conditional Layer Normalization (CondLN) and dual attention (Dual-attn) with and without CondLN. Single-attn uses only a self-attention block where the noised sequence, diffusion step, and history embeddings are directly combined. In contrast, Dual-attn integrates both self-attention and cross-attention layers to capture interactions between the noised sequence and user history. CondLN conditions the LN process on the noised sequence embedding $S_{t}$, diffusion time embedding $t$ adapting to personalized user behaviors. The results show that both CondLN and dual attention effectively boost performance with highest HR@5 performance when combining them. Meanwhile, the absence of any of these modules results in a decline in performance.
The effectiveness of DCRec. To verify the rationality of incorporating both implicit conditional and explicit conditional effectiveness in DCRec, we choose implicit conditional components only, explicit conditional components only (broadcasting the target item to the same length as historical interactions), and the original dual conditional diffusion transformer. As shown in Table 2, DCRec  consistently outperforms both ICDM and ECDM, demonstrating that jointly modeling implicit and explicit historical conditions is more effective than relying on either one individually.
Table 2. Performance comparison among Implicit Conditional DMs (ICDM), Explicit Conditional DMs (ECDM), and our DCDT. means the improvement is significant at $p<0.05$.
Dataset Metrics ICDM ECDM DCRec
Beauty HR@5 0.0553 0.0599 $0.0630^{*}$
HR@10 0.0839 0.0851 $0.0859^{*}$
NDCG@5 0.0376 0.0433 $0.0449^{*}$
NDCG@10 0.0482 0.0514 $0.0523^{*}$
Toys HR@5 0.0616 0.0633 $0.0690^{*}$
HR@10 0.0776 0.0803 $0.0903^{*}$
NDCG@5 0.0473 0.0500 $0.0518^{*}$
NDCG@10 0.0524 0.0571 $0.0587^{*}$
Yelp HR@5 0.0321 0.0349 $0.0405^{*}$
HR@10 0.0508 0.0551 $0.0611^{*}$
NDCG@5 0.0217 0.0241 $0.0272^{*}$
NDCG@10 0.0276 0.0320 $0.0338^{*}$
5.3.2. Hyper-parameters Analysis.
For a more nuanced understanding, we select certain sensitive hyper-parameter $\lambda$, and explore three different $\lambda$ settings as shown in Figure 6.
Figure 6. Figure (a) depicts three types of $\lambda$ and Figure (b) reports the corresponding HR@5 results on Amazon Beauty and Toys datasets. The $\lambda_{smooth}$ reports superior performance.
The loss balance factor $\lambda(\cdot)$ influences DCRec  performance by mediating the focus between recommendation and reconstruction tasks. While an increase in $\lambda$ initially bolsters performance by prioritizing embedding denoising, too high a value risks neglecting the core recommendation task, undermining overall performance. $\lambda_{fix}$ denotes experimental constant, $\lambda_{decline}$ denotes handcrafted function, $\lambda_{smooth}$ denotes training epoch related function, respectively. We can find that $\lambda_{smooth}$ can achieve the best performance. Due to the page limit, the details of $\lambda$ and other hyper-parameters please refer to Appendix E.
5.4. Model Efficiency Analysis (RQ3)
Figure 7. Effect of inference step $T$ of our DCRec  on Yelp and Beauty datasets ($T$ denotes the diffusion steps in training while $T^{\prime}$ denotes the steps in inference).
Figure 8. The impact of sample steps on our DCRec  and DiffuRec on HR@5 and time cost (second) on the Beauty dataset.
In this section, we investigate the impact of inference step $T$ to answer RQ3. We vary diffusion steps $T$ from 0 to $T$ in both Beauty and Toys datasets during inference and show the results in Figure 7. Moreover, we conducted a time comparison with DiffuRec (Li et al., 2024b), the most representative method in the baseline. To the fair comparison, we set 30 diffusion steps for training our DCRec  aligning to the DiffuRec as shown in Figure 8.
Firstly, We can find the best inference steps are vary in different datasets. This implies that the optimal diffusion step for each dataset is related to their item count. Then we observe that our DCDT ($f_{\theta}$) is optimized powerful enough to fulfill Equation 36 at an early $\hat{t}$-th stage, which significantly reduces time overhead. Compared to DiffuRec, our DCRec’s performance is more stable under different inference steps $T$ settings. In conclusion, our DCRec  can reach superior model performance than DiffuRec while maintaining lower inference time cost under different inference step settings. This is because our improvement enables the dual conditional mechanism, resulting in both accurate and efficient inference.
6. Conclusion
In this paper, we proposed the DCRec  to address key limitations in existing diffusion-based sequential recommender systems. By introducing a Markov chain to model item transitions and a dual conditional module to integrate contextual information, our approach improves the consistency of item selection and enhances the generation of target item embedding. Extensive experiments on public datasets validate the effectiveness of DCRec, demonstrating its superiority over state-of-the-art methods in both recommendation accuracy and computational efficiency.
Appendix A Algorithm of DCRec  Training and Inference
Algorithm 1 DCRec  Training
1:User interacted item embedding set $\{H_{0}\}$, item pool $\mathbb{Z}$, number of diffusion steps $T$, denoising network $f_{\theta}$, balance factor $\lambda$
2:repeat
3:     Sample ${H_{0}}$ from $\{H_{0}\}$ and corresponding $z$ from $\mathbb{Z}$
4:     $S_{0}$leftarrow Concatenate $H_{0}$ and $z$
5:     Sample $t\sim\mathcal{U}(1,T)$
6:     $S_{t}$leftarrow Adding noise to $S_{0}$ according to Equation 23
7:     Obtain predicted $\hat{S}_{0}\leftarrow f_{\theta}(S_{t},t,H_{0})$
8:     Extract predicted $\hat{e}_{0}$ from $\hat{S}_{0}$
9:     Obtain predicted $\hat{z}\leftarrow p_{\phi}(\hat{z}|\hat{e}_{0})$
10:     $\mathcal{L}_{T}$leftarrow $\|\hat{S}_{0}\|^{2}$
11:     $\mathcal{L}_{t}$leftarrow $\|\hat{e}_{0}-e_{0}\|^{2}$
12:     $\mathcal{L}_{z}$leftarrow $D_{\text{KL}}(q_{\phi}(z)||p_{\phi}(\hat{z}))$
13:     Take gradient descent step on$\nabla\left(\mathcal{L}_{T}+(1-\lambda)\mathcal{L}_{t}+\lambda\mathcal{L}_{z}\right)$ to optimize $\theta$ and $\phi$
14:until converged
15:optimized $\theta$, $\phi$.
Algorithm 2 DCRec  Inference
1:User interacted item embedding $H_{0}$, implicit conditioning scale factor $\delta$, number of diffusion steps $T$, number of accelerating steps $T^{\prime}$, denoising network $f_{\theta}$
2:Uniformly sample $T^{\prime}$ steps from $1$ to $T$
3:Sample $e_{T}~{}\sim\mathcal{N}(\bm{0},\bm{I})$
4:Sample $H_{T}~{}\sim\mathcal{N}(\bm{0},\bm{I})$
5:Implicit conditioning $H_{T}\leftarrow H_{T}+\delta H_{0}$
6:$S_{T}\leftarrow$ Concatenate $H_{T}$ and $e_{T}$
7:for $t\leftarrow T^{\prime}\dots 1$  do
8:     Obtain predicted ${S}_{0}\leftarrow f_{\theta}(S_{t},t,H_{0})$
9:     Transit to $S_{t-1}\leftarrow q(S_{t-1}|{S}_{0})$
10:Extract $e_{0}$ from $S_{0}$
11:obtain $z\leftarrow p_{\phi}(z|e_{0})$
12:$z$
Appendix B Dataset Preparation
We evaluate our proposed DCRec  on three publicly accessible datasets in different experiment settings. Table 3 provides a summary of the statistics for each dataset.
  • Amazon Beauty & Amazon Toys: We choose three representative sub-datasets from Amazon datasets: Beauty and Toys and keep the ‘5-core’ datasets (Zhou et al., 2020a) (Rendle et al., 2010)(Chen et al., 2022a), which filter out user-item interaction sequences with length less than 5.
  • Yelp: Yelp is a well-known business recommendation dataset. Following (Xie et al., 2022; Chen et al., 2022b), we only retain the transaction records after Jan. 1st, 2019 in our experiment.
For reproducibility, we follow the commonly used benchmark setting of RecBole (Zhao et al., 2021) to set up our experiments. Specifically, for each user, we first discard duplicated interaction sequences and sort the items in each user’s sequence chronologically by their timestamp. Furthermore, the maximum length of interaction sequences is set to 50. If there are more than 50 interactions in a sequence, we adopt the latest 50 interactions. If the number of interactions in a sequence is less than 50, we make it up to 50 by padding virtual items with the ID of 0. Following the common practice in sequential recommendations, we leave the interactions happening at the latest time as the test dataset and the interactions at the second latest time as the validation dataset (Chen et al., 2022a; Zhou et al., 2020b).
Table 3. Statistics of the datasets after preprocessing.
Dataset Beauty Toys Yelp
# Users 22,363 19,412 30,499
# Items 12,101 11,924 20,068
# Avg. Actions / User 8.9 8.6 10.4
# Avg. Actions / Item 16.4 14.1 15.8
# Actions 198,502 167,597 317,182
Sparsity 99.93% 99.93% 99.95%
Appendix C Baselines for Comparison
To evaluate the performance of our model DCRec, we select various representative and/or state-of-the-art recommendation models as baselines.
  • GRU4Rec (Hidasi and Karatzoglou, 2018): This model uses the GRU to model the user interaction sequence and gives the final recommendation.
  • Caser (Yuan et al., 2019): CNN-based approach for SR.
  • SARSRec (Kang and McAuley, 2018): This model employs a single-direction Transformer with a masked encoder to capture explicit correlations between items.
  • Bert4Rec (Sun et al., 2019): replaces the next-item prediction with a Cloze task to fuse information between an item (a view) in a user behavior sequence and its contextual information.
  • $S^{3}$ec (Zhou et al., 2020c): uses SSL to capture correlation-ship among items, sub-sequences, and associated attributes from the given user behavior sequence. Its modules for mining on attributes are removed because we don’t have attributes for items.
  • CL4SRec (Xie et al., 2022): fuses contrastive SSL with a Transformer-based SR model.
  • ICLRec (Chen et al., 2022b): leverages latent user intents, learned from unlabeled behavior sequences, to optimize SR (SR) models using contrastive self-supervised learning.
  • DreamRec (Yang et al., 2023b): uses the historical interaction sequence as conditional guiding information for the diffusion model to enable personalized recommendations.
  • SdifRec (Xie et al., 2024): introduce the Schrödinger Bridge into diffusion-based SR, by replacing the Gaussian prior with the user’s current state.
  • DiffuRec (Li et al., 2024b): incorporates a diffusion model into the field of SR, leveraging a Transformer backbone to reconstruct target item representations based on the user’s historical interaction behaviors.
Appendix D Ablation Study for Partially Noised Input Sequence
In this paper, we concatenate the history embedding and target item embedding as a sequential input for our denoising model DCDT. As shown in Figure 2, we add noise to the whole input sequence. In this section, motivated by DiffuSeq (Gong et al., 2023) which is another diffusion model for Q&A task in NLP, we analyze the effectiveness of only adding noise to the target item. The results are shown in Figure 9. It seems that partially adding noise to the target item is worse than adding noise to the whole sequence. The possible reason is the unbalanced noised sequence length. DiffuSeq concatenates the question and answer as an input and only adds noise to the answer sequence. In that case, the question sequence and answer sequence have similar sequence lengths. However, in our SR task, the history sequence is nearly 50×50$\times$50 × longer than the target. Therefore, the strategy of partially noised input is unfeasible for DCRec.
Figure 9. Performance comparison of history conditioning with noise and without noise.
Appendix E Details of Hyper-parameters Analysis
E.1. The Balance Factor $\lambda$
It’s worth to note that the optimization in Equations 31 and 24 is a complex and parameter-sensitive multi-objective task. In this section, we detailed analyze the impacts of three types of different $\lambda$ settings, including fix factor $\lambda_{\text{fix}}$, decline factor $\lambda_{\text{decline}}$ and smoothly decline factor $\lambda_{\text{smooth}}$. We formulate them as follows:
$\lambda_{\text{fix}}&=c\\ \lambda_{\text{decline}}&=\text{Max}-g_{d}(\text{ep})+\text{Min}\\ \lambda_{\text{smooth}}&=\text{Max}*g_{s}(\text{ep})+\text{Min}\\ \end{aligned}\right.$
where $c$ is a constant and ep is the current training epoch. Max and Min denote the upper bound and the lower bound of the decline factor $\lambda_{\text{decline}}$ and $\lambda_{\text{smooth}}$. $g_{d}(\text{ep})$ is a decline linear function associated with the current training epoch, while $g_{s}(\text{ep})$ is another decline exponential function also associated with the epoch. In Table 4 and Figure 10, we illustrate several example settings of $\lambda$ in our experiments and corresponding results are shown in Table 5. We note that those declining settings dramatically improve the model performance compared to the fixed ones. The $\lambda_{\text{smooth}}$ settings are feasible for Toys and Beauty datasets, while $\lambda_{\text{decline}}$ settings are more suitable for the Yelp dataset.
Table 4. Detailed settings of different $\bm{\lambda}$.
Tag Type Function
$\lambda_{\text{f1}}$ $\lambda_{\text{fix}}$ 0.2
$\lambda_{\text{f2}}$ $\lambda_{\text{fix}}$ 0.1
$\lambda_{\text{f3}}$ $\lambda_{\text{fix}}$ 0.003
$\lambda_{\text{d1}}$ $\lambda_{\text{decline}}$$\begin{cases}0.2-0.2*\text{ep}/20+3\times 20^{-3}&0<\text{ep}<20\\ 3\times 10^{-3}&\text{ep}>=20\\ \end{cases}$
$\lambda_{\text{d2}}$ $\lambda_{\text{decline}}$$\begin{cases}0.2-0.2*\text{ep}/30+3\times 10^{-3}&0<\text{ep}<30\\ 3\times 10^{-3}&\text{ep}>=30\\ \end{cases}$
$\lambda_{\text{d3}}$ $\lambda_{\text{decline}}$$\begin{cases}0.2-0.2*\text{ep}/40+3\times 10^{-3}&0<\text{ep}<40\\ 3\times 10^{-3}&\text{ep}>=40\\ \end{cases}$
$\lambda_{\text{d4}}$ $\lambda_{\text{decline}}$$\begin{cases}0.2-0.2*\text{ep}/80+3\times 10^{-3}&0<\text{ep}<80\\ 3\times 10^{-3}&\text{ep}>=80\\ \end{cases}$
$\lambda_{\text{s1}}$ $\lambda_{\text{smooth}}$$\begin{cases}2*(1-\text{ep}/150)^{6}+3\times 10^{-3}&0<\text{ep}<150\\ 3\times 10^{-3}&\text{ep}>=150\\ \end{cases}$
$\lambda_{\text{s2}}$ $\lambda_{\text{smooth}}$$\begin{cases}2*(1-\text{ep}/150)^{9}+3\times 10^{-3}&0<\text{ep}<150\\ 3\times 10^{-3}&\text{ep}>=150\\ \end{cases}$
$\lambda_{\text{s3}}$ $\lambda_{\text{smooth}}$$\begin{cases}0.05*(1-\text{ep}/150)^{10}+3\times 10^{-3}&0<\text{ep}<150\\ 3\times 10^{-3}&\text{ep}>=150\\ \end{cases}$
Figure 10. Illustrations of different $\bm{\lambda}$.
Table 5. The comparison performance of different $\bm{\lambda}$ settings. Bold scores represent the highest results.
Dataset Metric $\lambda_{\text{fix}}$ $\lambda_{\text{decline}}$ $\lambda_{\text{smooth}}$
$\lambda_{f1}$ $\lambda_{f2}$ $\lambda_{f3}$ $\lambda_{d1}$ $\lambda_{d2}$ $\lambda_{d3}$ $\lambda_{d4}$ $\lambda_{s1}$ $\lambda_{s2}$ $\lambda_{s3}$
Beauty HR@5 0.0453 0.0507 0.0493 0.0320 0.0426 0.0479 0.0584 0.0617 0.0630 0.0564
HR@10 0.0643 0.0704 0.0669 0.0471 0.0636 0.0688 0.0844 0.0855 0.0859 0.0795
NDCG@5 0.0318 0.0366 0.0357 0.0224 0.0298 0.0340 0.0414 0.0435 0.0449 0.0399
NDCG@10 0.0379 0.0429 0.0413 0.0272 0.0365 0.4065 0.0498 0.0514 0.0522 0.0473
Toys HR@5 0.0470 0.0582 0.0458 0.0113 0.0284 0.0396 0.0676 0.0684 0.0690 0.0594
HR@10 0.0627 0.0787 0.0648 0.0195 0.0418 0.0552 0.0889 0.0901 0.0903 0.0804
NDCG@5 0.0352 0.0436 0.0329 0.0083 0.0203 0.0287 0.0503 0.0510 0.0518 0.0447
NDCG@10 0.0402 0.0502 0.0390 0.0109 0.0244 0.0337 0.0571 0.0579 0.0587 0.0514
Yelp HR@5 0.0241 0.0269 0.0247 0.0344 0.0405 0.0396 0.0251 0.0327 0.0289 0.0398
HR@10 0.0363 0.0411 0.0424 0.0547 0.0611 0.0605 0.0174 0.0514 0.0192 0.0591
NDCG@5 0.0169 0.0179 0.0163 0.0232 0.0272 0.0266 0.0388 0.0223 0.0443 0.0270
NDCG@10 0.0207 0.0225 0.0220 0.0297 0.0338 0.0333 0.0218 0.0282 0.0241 0.0333
E.2. Other Hyper-parameters Analysis
Figure 11. Hyperparameters analysis of our DCRec  on Beauty dataset.
We report the experimental results for key hyperparameters on the Beauty dataset, including the noise schedule $\beta$, number of diffusion steps $T$, and scaling factor $\delta$. In detail, $\beta$ is tuned within $\{0.01,0.02,0.04,0.06,0.08,0.1\}$, diffusion steps $T$ is tuned within $\{5,10,15,20,30,50,75,100,150,200\}$ and scaling factor $\delta$ is tuned within $\{0.1,0.2,0.5,0.75,1\}$. As shown in Figure 11, we can find for the noise schedule $\beta$, performance remains stable for values up to 0.04 but sharply declines at 0.06, indicating that excessive noise negatively impacts the model. The number of training diffusion steps $T$ improves performance up to 20 steps, after which it plateaus, suggesting that 20 steps are sufficient. The scaling factor $\delta$ performs best in the range of 0.2 to 0.5, with both smaller and larger values resulting in decreased performance.
Appendix F Implementation Details
We implement our model and baselines using Pytorch (Paszke et al., 2019) and conduct our experiments on an NVIDIA RTX 3090 GPU with 24G memory. For a fair comparison, we carefully tuned the model parameters of all the baselines on the validation set. Then we choose the model parameters that achieve the best performance on the validation set to compare. For our DCRec, based on the hyperparameter analysis, we set the batch size to 512, the item embedding dimension to 128, learning rate of $34\times 1e-$ 3, maximum sequence length to 50, $\beta$ to 0.02, the dropout rate to 0.1, the embedding dropout to 0.4, the hidden factor to 128, $\delta$ to 0.1, the number of transformer blocks to 4, and the number of diffusion steps $T$ to 50. For $\lambda$, we adopt $\lambda_{smooth}=\lambda_{s2}$ for the Beauty and Toys datasets, while for the Yelp dataset, we use $\lambda_{decline}=\lambda_{d2}$, as determined by the results in Table 5 result. We choose Adam (Kingma and Ba, 2015) as our optimization strategy and use early stopping with a patience of 5 to prevent overfitting.

### v2

@@
 Recent advancements in diffusion models have shown promising results in sequential recommendation (SR).
-However, current diffusion-based methods still exhibit two key limitations.
-First, they implicitly model the diffusion process for target item embeddings rather than the discrete target item itself, leading to inconsistency in the recommendation process.
-Second, existing methods rely on either implicit or explicit conditional diffusion models, limiting their ability to fully capture the context of user behavior and leading to less robust target item embeddings.
-In this paper, we propose the Dual Conditional Diffusion Models for Sequential Recommendation (DCRec), introducing a discrete-to-continuous sequential recommendation diffusion framework.
-Our framework introduces a complete Markov chain to model the transition from the reversed target item representation to the discrete item index, bridging the discrete and continuous item spaces for diffusion models and ensuring consistency with the diffusion framework.
-Building on this framework, we present the Dual Conditional Diffusion Transformer (DCDT) that incorporates the implicit conditional and the explicit conditional for diffusion-based SR.
-Extensive experiments on public benchmark datasets demonstrate that DCRec  outperforms state-of-the-art methods.
+Existing approaches predominantly rely on implicit conditional diffusion models, which compress user behaviors into a single representation during the forward diffusion process.
+While effective to some extent, this oversimplification often leads to the loss of sequential and contextual information, which is critical for understanding user behavior.
+Moreover, explicit information, such as user-item interactions or sequential patterns, remains underutilized, despite its potential to directly guide the recommendation process and improve precision.
+However, combining implicit and explicit information is non-trivial, as it requires dynamically integrating these complementary signals while avoiding noise and irrelevant patterns within user behaviors.
+To address these challenges, we propose Dual Conditional Diffusion Models for Sequential Recommendation (DCRec), which effectively integrates implicit and explicit information by embedding dual conditions into both the forward and reverse diffusion processes.
+This allows the model to retain valuable sequential and contextual information while leveraging explicit user-item interactions to guide the recommendation process.
+Specifically, we introduce the Dual Conditional Diffusion Transformer (DCDT), which employs a cross-attention mechanism to dynamically integrate explicit signals throughout the diffusion stages, ensuring contextual understanding and minimizing the influence of irrelevant patterns.
+This design enables precise and contextually relevant recommendations.
+Extensive experiments on public benchmark datasets demonstrate that DCDRec significantly outperforms state-of-the-art methods in both accuracy and computational efficiency.
 Equal contribution.
-copyright: nonecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
+copyright: nonecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
 1. Introduction
 Sequential Recommender Systems (SRSs) are designed to produce recommendations by predicting the next item that will capture a user’s interest (Wang et al., 2019a).
-Despite their effectiveness, existing models represent items with fixed vectors, which limit their ability to capture the latent aspects of items as well as the diverse and uncertain nature of user preferences (Li et al., 2024b; Yang et al., 2023b).
-Besides, these methods assume that the items with the most user interactions are the most relevant, which puts less-interacted items at a disadvantage and leads to the exposure problem (Yang et al., 2023b).
-Diffusion models (Ho et al., 2020; Ho and Salimans, 2022a), as an emergent deep generative model, add random noise to input data in a forward process and then recover the original data through a step-by-step reverse denoising process.
-Their denoising nature aligns well with the SR task, as modeling sequential item recommendations mirrors this step-wise approach (Xie et al., 2024).
-Recent studies on diffusion models for SR (Yang et al., 2023b) follow the paradigm of diffusion models, perturbing the embedding of the target item through a forward diffusion process into a Gaussian prior distribution.
+Despite effectiveness, existing models represent items with fixed vectors, which limit their ability to capture the latent aspects of items as well as the diverse and uncertain nature of user preferences (Li et al., 2024b; Yang et al., 2023b).
+Besides, these methods assume that the items with the most user interactions are the most relevant, which puts less-interacted items at a disadvantage and leads to the exposure problem (Yang et al., 2023b, 2024a).
+Diffusion models (Ho et al., 2020; Ho and Salimans, 2022a), add random noise to input data in a forward process and then recover the original data through a step-by-step reverse denoising process.
+Their denoising nature aligns well with the SR task, as this step-wise approach can mimic the modeling sequential item recommendations (Xie et al., 2024).
+Recent studies on diffusion models for SR (Yang et al., 2023b; Li et al., 2024b) follow the paradigm of diffusion models, perturbing the embedding of the target item through a forward process into a Gaussian prior distribution.
 Then, they restore the Gaussian distribution iteratively through a reverse denoising process, also referred to as the sampling phase, to recover target representations and recommend items that are most similar to it.
-Although these methods have achieved significant performance, current diffusion models for SR are still undergoing some shortcomings.
-Firstly, existing approaches (Yang et al., 2023b) commonly conduct conditional diffusion-based recommendations by conditioning the reverse denoising process on user behavior sequences.
-These methods focus on optimizing the diffusion process for the target item embeddings or generating an oracle item embedding for recommendation (Lin et al., 2024b).
-However, they do not explicitly model the diffusion process for the discrete target item itself (Gap 1).
-As illustrated in Figure 1(a), the traditional methods overlook a critical step: mapping the reversed target item representation into the discrete item index space.
-These methods often determine the recommended item by calculating the similarity (e.g., inner product) between the reversed target item representation and candidate item embeddings, selecting the item with the highest similarity score.
-While this method works for item ranking, it does not align with the core principles of diffusion models.
-Diffusion models are inherently probabilistic and continuous, but this step—based on a direct inner product comparison—introduces a deterministic, discrete decision process that breaks the continuous generative flow of diffusion.
-This inconsistency creates a gap between the optimization directions of the ranking loss for the recommendation task and the denoising loss for the diffusion model, resulting in suboptimal recommendation performance.
-Figure 1.
-Figures (a) and (b) illustrate the differences between previous methods and our proposed approach in Markov modeling, while Figures (c) and (d) highlight the contrast between implicit and explicit conditional diffusion models for SR.
-More importantly, existing diffusion-based SR methods can be broadly classified into two categories: implicit and explicit conditional approaches.
-As shown in Figure 1(c) and (d), in this paper, Implicit conditional methods refer to treating the user’s historical behavior as an implicit feature and integrating them into the generating target item at each diffusion step (Yang et al., 2023b; Li et al., 2024a, b).
-In contrast, Explicit conditional methods refer to leverage those historical behaviors as explicit guidance that directly influences the denoising process in diffusion SR (Lin et al., 2024a).
-The main difference between the two types of conditioning is that the implicit one can be regarded as a feature fusion with fixed and compressed historical features while the explicit one iteratively interacts with the generating target according to the diffusion progress.
-However, both approaches have inherent limitations.
-Implicit conditioning may oversimplify complex user behaviors by compressing them into a single vector, potentially losing valuable sequential information.
-Explicit conditioning, while providing detailed conditioning, can be more sensitive to noise and irrelevant actions within the user’s behavior sequence, which may hinder the recommendation quality (He and McAuley, 2016a).
-As no existing work combines implicit and explicit conditioning to leverage their respective strengths for enhanced SR (Gap 2).
+Diffusion-based sequential recommendation differs from standard diffusion in computer vision tasks, as recommender systems aim to fulfill user interests.
+The generated target should elicit positive user feedback, making it essential for the diffusion model to be more controllable with respect to user preferences.
+Figure 1. Figures (a) and (b) are examples of implicit and explicit conditional diffusion models for SR.
+In this paper, we broadly classify diffusion-based SR into two categories: implicit and explicit conditional approaches, as shown in Figure 1(a) and (b).
+Implicit conditional methods refer to treating the user’s historical behavior as an implicit feature and integrating them into the generating target item at each diffusion step (Yang et al., 2023b; Li et al., 2024a, b; Xie et al., 2024).
+In contrast, Explicit conditional methods refer to leveraging those historical behaviors as explicit guidance that directly influences the denoising process in diffusion SR.
+The main difference between these two types of conditioning is that the implicit one can be regarded as feature fusion with historical features while the explicit one iteratively interacts with the generating target according to the diffusion progress.
+Currently, most diffusion-based SR methods are implicit conditional approaches (Yang et al., 2023b; Li et al., 2024a, b; Xie et al., 2024).
+While promising to some extent, implicit conditioning may oversimplify complex user behaviors by compressing them into a single vector, potentially losing valuable sequential patterns and contextual information.
+Additionally, explicit conditioning, which offers more precise signals, has shown its potential for directly guide the recommendation process in reranking tasks (Lin et al., 2024a), though it remains unexplored in the sequential recommendation.
+Moreover, there is no existing work that combines implicit and explicit conditioning to leverage their respective strengths, which is non-trivial for providing more precise recommendations based on user preference..
+It requires dynamically integrating these complementary signals to fulfill the diffusion optimization paradigm while avoiding noise and irrelevant patterns within user behaviors.
 Another problem for existing diffusion models-based SR is their inefficient and slow sampling process, which relies on long Markov chains.
 Generating a target item requires iteratively refining random noise through numerous denoising steps, making the process computationally intensive and time-consuming (Lin et al., 2024b).
-Aiming at bridging the aforementioned two significant gaps in existing SRSs, we propose the Dual Conditional Diffusion Models for SR (DCRec).
-Firstly, we begin with introducing the discrete-to-continuous sequential diffusion framework for SR.
-Our approach maintains consistency with the existing diffusion framework by introducing a Markov chain to model the transition from the reversed target item representation to the discrete item index.
-This probabilistic step fits more naturally within the diffusion process, allowing us to better capture the stochastic nature of item generation and selection.
-By explicitly modeling this transition, our method not only aligns with the underlying diffusion model but also provides a more coherent and accurate recommendation process.
-Furthermore, instead of following the existing implicit conditional or explicit conditional diffusion model-based SR methods, we propose novel conditioned diffusion models that integrate both implicit and explicit conditioning strategies.
+We propose the Dual Conditional Diffusion Model for Sequential Recommendation (DCRec).
+Specifically, instead of employing a single condition, we propose a novel conditioned diffusion model framework that integrates both implicit and explicit conditioning strategies.
+Such an approach comprehensively accounts for both explicit and implicit advantages while preserving the evidence lower bound (ELBO) consistency under the diffusion framework.
 Based on our sequential recommendation diffusion framework, we introduce the Dual Conditional Diffusion Transformer (DCDT).
-This model concatenates historical information as an implicit condition while leveraging a cross-attention mechanism to incorporate explicit control signals throughout the diffusion process.
+This module concatenates historical information and the expected target item as an implicit condition while leveraging a cross-attention mechanism to incorporate explicit control signals dynamically throughout the diffusion process.
 By embedding dual conditions into both the forward and reverse diffusion stages, our approach ensures more precise and contextually relevant recommendations.
 In addition, the dual conditional mechanism leads to improved model performance, allowing the inference process to achieve optimal results with only a few sampling steps.
-This reduction in sampling steps significantly lowers computational overhead, making our model more efficient and better suited for real-time applications.
+This reduction in sampling steps significantly lowers computational overhead, making our model more efficient and better suited for online applications.
 The main contributions of this work are threefold:
-  • We propose the Dual Conditional Diffusion Models for the sequential recommendation, introducing a novel diffusion framework for the SR.
-We explicitly introduce the discrete target items into the continuous diffusion process, defining the unified optimization target that considers both the ranking and the diffusion task.
-  • Based on this framework, we propose a unified module, named Dual Conditional Diffusion Transformer (DCDT), which can generate accurate target item embedding based on both implicit and explicit conditions.
-Furthermore, our framework simplifies the diffusion process by enabling few-step inference instead of the traditional full-step inference.
-  • We conduct extensive experiments on various public benchmark datasets and show that DCRec  outperforms the state-of-the-art SR models.
-2. Related Work
-$\bullet$ Diffusion Models.
-Diffusion models, derived from non-equilibrium thermodynamics, have achieved notable success in areas like computer vision, sequence modeling, and audio processing (Yang et al., 2024).
-Two primary paradigms exist unconditional models and conditional models.
-Unconditional models, such as DDPM (Ho et al., 2020), aim to push the performance boundaries of generative models.
-Conditional models incorporate guidance from labels to control generated samples.
-SGGM (Yang et al., 2022) conditions on self-produced labels, while You et al.
-(You et al., 2023) demonstrate how few labels and pseudo-training enhance both large-scale diffusion models and semi-supervised learners.
-Dhariwal (Dhariwal and Nichol, 2021) uses classifier guidance to improve sample quality, and Ho and Salimans (Ho and Salimans, 2022b) combine conditional and unconditional scores to balance sample quality and diversity.
-Multi-modal guidance further enriches the diffusion process, as seen in DiffuSeq (Gong et al., 2023) for NLP tasks and SDEdit (Meng et al., 2022) for image translation, with latent diffusion models (LDM) (Rombach et al., 2022) offering flexible, unified semantics.
-unCLIP (Ramesh et al., 2022) and ConPreDiff (Yang et al., 2023a) integrate CLIP latents in text-to-image generation.
-$\bullet$ Diffusion Model-based Sequential Recommendation.
-Diffusion models show promise in the sequential recommendation by using user behavior sequences as guidance in the denoising process (Xie et al., 2024; Yang et al., 2023b; Li et al., 2024b).
-DiffuRec (Li et al., 2024b) corrupts target item embeddings with Gaussian noise and uses a Transformer to iteratively reconstruct them.
-DreamRec (Yang et al., 2023b) generates the oracle next-item embeddings based on user preferences but struggles with scalability due to its lack of negative sampling.
-DCDR (Lin et al., 2024a) adopts a discrete diffusion framework for progressive re-ranking.
-However, current methods face two main issues: they focus on embedding diffusion rather than discrete item generation, leading to inconsistencies, and they rely on implicit or explicit conditioning, limiting the capture of user context and weakening target item embeddings.
+  • We first propose Dual Conditional Diffusion Models for the sequential recommendation, introducing a novel diffusion framework for SR.
+  • Based on this framework, we propose a unified module, Dual Conditional Diffusion Transformer (DCDT), which can generate accurate target item embeddings based on both implicit and explicit conditions.
+Furthermore, our framework accelerates the diffusion inference process by significantly reducing the sampling steps.
+  • We conduct extensive experiments on various public benchmark datasets and show that DCRec  outperforms state-of-the-art SR models.
+Comprehensive empirical analyses and ablation studies further demonstrate the effectiveness of the proposed framework.
 2. Preliminary
+Before presenting our model, we define the scope of the problem in sequential recommendation.
+Subsequently, we briefly outline the theoretical foundation of vanilla diffusion models for continuous spaces and current diffusion-based sequential recommender models for discrete spaces.
 2.1. Problem Formulation
@@
 Generally, for a user $u$, given their sequence context $H_{z}=\{z_{1},z_{2},...,z_{n-1}\}$, the task of an SRS is to predict the next item (target item) $z$ which may interest $u$.
-3.2. Diffusion Models for Continuous Domains
-Recently, diffusion models have already demonstrated significant performance in continuous domains like computer vision and audio (Ho et al., 2020; Ho and Salimans, 2022a; Kong et al., 2021; Rombach et al., 2022).
-Thus, before presenting our diffusion SR model, we briefly recap the diffusion models in continuous domains (Ho et al., 2020; Luo, 2022).
-A generative model is expected to match estimation distribution $p_{\theta}(e_{0})$ to the ground truth distribution $p(e_{0})$ as closely as possible, where $\theta$ represents learnable parameters and $e_{0}\in\mathbb{R}^{d}$ is a set of samples.
-For vanilla continuous diffusion models, they model the sample $e_{0}$ as a Markov chain with a sequence of latent variables $\{e_{T},\dots,e_{t},e_{t-1},\dots,e_{0}\}$, where $T$ is the diffusion step and $e_{T}$ is a standard Gaussian, $e_{t}$ is the intermediate latent variable.
-By this means, the latent dimension is exactly equal to the data dimension of $e_{0}$, and the diffusion model progressively recovers $e_{0}$ from Gaussian noise.
-Each denoising transition $e_{t}\rightarrow e_{t-1}$ is parameterized by a linear Gaussian transition, calculated by a denoising DNN $f_{\theta}(e_{t},t)$:
-(1) $p_{\theta}(e_{t-1}|e_{t})=\mathcal{N}(e_{t-1};\mu_{\theta}(e_{t},t),\Sigma_{\theta}(e_{t},t))$
-Thus, the joint distribution of diffusion models can be written as:
-(2) $\small p_{\theta}(e_{0:T})=p_{\theta}(e_{T})\prod^{T}_{t=1}p_{\theta}(e_{t-1}|e_{t})$
-(3) $where,p_{\theta}(e_{T})=\mathcal{N}(e_{T};\mathbf{0},\mathbf{I})$
-To train the denoising model $f_{\theta}(e_{t},t)$, diffusion models first construct a progressive noising process $\{e_{0},\dots,e_{T-1},e_{T}\}$, known as the forward process.
-Corresponding to this, the denoising process is called the reverse process.
-The forward process adds Gaussian noise to $e_{0}$ step-by-step until the final latent $e_{T}$ is guaranteed to a standard Gaussian.
-Each noising transition $e_{t-1}\rightarrow e_{t}$ is defined as a linear Gaussian transition $q(e_{t}|e_{t-1})=\mathcal{N}(e_{t};\sqrt{\alpha_{t}}e_{t-1},(1-\alpha_{t})\mathbf{I})$, where the Gaussian parameters $\alpha_{t}\in(0,1)$ varies over time step $t$ and controls the noise scale added to the embedding in the current step $t$.
-Due to the Markov property and the Bayes’ theorem, the transition can be rewritten with $e_{0}$ as:
-(4) $q(e_{t}|e_{t-1})=q(e_{t}|e_{t-1},e_{0})=\frac{q(e_{t-1}|e_{t},e_{0})q(e_{t}|e_{0})}{q(e_{t-1}|e_{0})}$
-To optimize the parameter $\theta$, the diffusion model is trained to maximize the log-likelihood of $p_{\theta}(x)$ according to observed samples $e_{0}$.
-This objective can be formulated as $\arg\max_{\theta}\log p_{\theta}(e_{0})$.
-Mathematically, we can derive a proxy objective called the Evidence Lower Bound (ELBO), a lower bound of the likelihood term $\log p_{\theta}(e_{0})$.
-Formally, the equation of the ELBO is:
-(5) $\mathbb{E}_{q(e_{1:T}|e_{0})}\left[\log\frac{p_{\theta}(e_{0:T})}{q(e_{1:T}|e_{0})}\right]$
-Rather than directly maximize the likelihood term, diffusion models minimize the minus ELBO as (Luo, 2022; Ho et al., 2020):
-(6) $\arg\max\limits_{\theta}\log p_{\theta}(e_{0})$ $\approx\arg\max\limits_{\theta}\mathbb{E}_{q(e_{1:T}|e_{0})}\left[\log\frac{p_{\theta}(e_{0:T})}{q(e_{1:T}|e_{0})}\right]$
-(7) $=\arg\min\limits_{\theta}\mathbb{E}_{q(e_{1:T}|e_{0})}\left[\log\frac{q(e_{1:T}|e_{0})}{p_{\theta}(e_{0:T})}\right]$
-Combining with Equations 2 and 4, the log term in Equation 7 can be further derived as:
-(8) $\log\frac{q(e_{T}|e_{0})}{p_{\theta}(e_{T})}+\sum_{t=2}^{T}\log\frac{q(e_{t-1}|e_{t},e_{0})}{p_{\theta}(e_{t-1}|e_{t}))}-\log p_{\theta}(e_{0}|e_{1})$
-Therefore, based on Equations 7 and 8, the ELBO-based optimization object of parameter $\theta$ is to minimize:
-(9) $\mathcal{L}_{\text{elbo}}(e_{0})=\mathbb{E}_{q(e_{1:T}|e_{0})}$ $\Big{[}\log\frac{q(e_{T}|e_{0})}{p_{\theta}(e_{T})}+\sum_{t=2}^{T}\log\frac{q(e_{t-1}|e_{t},e_{0})}{p_{\theta}(e_{t-1}|e_{t})}$
-$-\log p_{\theta}(e_{0}|e_{1})\Big{]}$
-To simplify and combine the second and third terms in Equation 9, recent research (Luo, 2022) derivative a simple surrogate objective to obtain a mean-squared error term:
-(10) $\mathcal{L}_{\text{elbo}}(e_{0})=\mathbb{E}_{q(e_{1:T}|e_{0})}\left[\sum_{t=1}^{T}\|f_{\theta}(e_{t},t)-e_{0}\|^{2}\right]$
-4. Dual Conditional Diffusion Models for Sequential Recommendation
 Figure 2.
 (a) illustrates a simplified workflow of our DCRec.
-During the end-to-end diffusion process, DCRec  implicitly concatenates the noised target item embedding $\bm{e_{t}}$ and the corresponding noised history embedding sequence $\bm{H_{t}}$ at each diffusion step.
+During the end-to-end diffusion process, DCRec  implicitly concatenates the noisy target item embedding $\bm{e_{t}}$ and the corresponding noisy history embedding sequence $\bm{H_{t}}$ at each diffusion step.
 Meanwhile, DCRec  is also guided by clear history embedding $\bm{H_{0}}$ as explicit conditional signals.
@@
 The dual conditional trajectory with additional explicit guidance is more likely to reach the target item (the orange area).
-In this section, we present the comprehensive introduction of our DCRec  as shown in Figure 2, which includes both the Sequential Recommendation Diffusion Framework and the Dual Conditional Diffusion Transformer (DCDT).
-Firstly, we expend the vanilla continuous diffusion to the discrete sequential recommendation task and provide a theoretical-grounded diffusion framework.
-Then, we introduce the DCDT, our denoising network $f_{\theta}$ for diffusion SR under the proposed framework.
-Finally, we focus on the efficient inference for the DCRec.
-4.1. Sequential Recommendation Diffusion Framework
-Continuous diffusion models have achieved remarkable success in domains like vision and audio (Kong et al., 2021), but their application to SR has been limited due to the inherently discrete nature of the recommended items.
-Existing diffusion models in SR typically determine the recommended item by computing the similarity (e.g., inner product) between the reversed target item representation and candidate item embeddings, selecting the item with the highest similarity score.
-While this approach is effective for ranking, it fundamentally conflicts with the principles of diffusion models.
-Vanilla diffusion models are designed to be probabilistic and continuous.
-Current diffusion SR approaches (Yang et al., 2023b; Li et al., 2024b; Xie et al., 2024) directly apply the continuous diffusion framework to the SR task neglecting the discrete-to-continuous transition.
-This breaks the transition Markov chain, the generative flow that diffusion models rely on, creating a disconnect between the ranking objective and the diffusion model’s denoising process.
-Specifically, this misalignment leads to an inconsistency between the optimization directions of the ranking loss for recommendation and the denoising loss of the diffusion process, resulting in sub-optimal recommendation performance.
-To address this issue and motivated by diffusion models in text domains (Li et al., 2022), we explicitly extend continuous diffusion models to discrete item domains as shown in Figure 2(a).
-As for discrete target item $z$ from the item pool $\mathbb{Z}$, the Markov chain in the forward and reverse processes are extending as $\{z,\dots,e_{T-1},e_{T}\}$ and $\{e_{T},\dots,e_{0},z\}$.
-Specifically, to map the discrete variables into continuous domains, we define a learnable embedding function $\text{Emb}(z)$.
-The forward transition process of $z$ is defined as:
+2.2. Diffusion Models
+In this section, we first recap the paradigm of vanilla diffusion models for continuous domains like computer vision and audio (Ho et al., 2020; Ho and Salimans, 2022a; Kong et al., 2021; Rombach et al., 2022).
+Then, we briefly introduce the expansion of diffusion models for discrete sequential recommendation task.
+2.2.1. Vanilla Diffusion Models
+Diffusion Models is expected to match estimation distribution $p_{\theta}(e_{0})$ to the ground truth distribution $p(e_{0})$ as closely as possible, where $\theta$ represents learnable parameters and $e_{0}\in\mathbb{R}^{d}$ is a set of samples.
+They recover the sample $e_{0}$ through a progressive denoising process modeled as a Markov chain $\{e_{T},e_{T-1},\dots,e_{0}\}$, where $T$ is the diffusion step and $e_{T}$ is a standard Gaussian, $e_{t}$ is the intermediate latent variable.
+Each denoising transition $e_{t}\rightarrow e_{t-1}$ is parameterized by a linear Gaussian transition.
+The arithmetic mean $\mu_{\theta}$ and standard deviation $\Sigma_{\theta}$ are calculated by a denoising DNN $f_{\theta}(e_{t},t)$.
+Therefore, the transition with trainable parameters $\theta$ can be formulated as follows:
+(1) $p_{\theta}(e_{t-1}|e_{t})=\mathcal{N}(e_{t-1};\mu_{\theta}(e_{t},t),\Sigma_{\theta}(e_{t},t)),$
+Diffusion models train the denoising model $f_{\theta}(e_{t},t)$ by constructing a progressive noising process $\{e_{0},\dots,e_{T-1},e_{T}\}$, known as the forward process, with the aforementioned denoising process referred to as the reverse process.
+The forward process adds Gaussian noise to $e_{0}$ step-by-step until the final latent $e_{T}$ is guaranteed to a standard Gaussian.
+Each noising transition $e_{t-1}\rightarrow e_{t}$ is defined as a linear Gaussian transition:
+(2) $q(e_{t}|e_{t-1})=\mathcal{N}(e_{t};\sqrt{\alpha_{t}}e_{t-1},(1-\alpha_{t})\mathbf{I}),$
+where the Gaussian parameters $\alpha_{t}\in(0,1)$ varies over time step $t$ and controls the noise scale added to the embedding in the current step $t$.
+The diffusion model is trained to maximize the log-likelihood of $p_{\theta}(x)$ according to observed samples $e_{0}$, which can be formulated as $\arg\max_{\theta}\log p_{\theta}(e_{0})$.
+Following (Ho et al., 2020), the canonical objective of a diffusion model is to optimize the Evidence Lower Bound (ELBO) of $\log p_{\theta}(e_{0})$:
+(3) $\mathcal{L}_{\text{elbo}}(e_{0})=\mathbb{E}_{q(e_{1:T}|e_{0})}$ $\Big{[}\log\frac{q(e_{T}|e_{0})}{p_{\theta}(e_{T})}+\sum_{t=2}^{T}\log\frac{q(e_{t-1}|e_{t},e_{0})}{p_{\theta}(e_{t-1}|e_{t})}$
+$-\log p_{\theta}(e_{0}|e_{1})\Big{]},$
+To simplify Equation 3, recent research (Luo, 2022) derivatives a simple surrogate objective to obtain a mean-squared error term:
+(4) $\mathcal{L}_{\text{elbo}}(e_{0})=\mathbb{E}_{q(e_{1:T}|e_{0})}\left[\sum_{t=1}^{T}\|f_{\theta}(e_{t},t)-e_{0}\|^{2}\right],$
+2.2.2. Diffusion Models for Sequential Recommendation
+The inherently discrete item domain in sequential recommendation poses challenges for applying standard diffusion models, which are originally designed for continuous domains.
+Considering discrete target item $z$ from the item pool $\mathbb{Z}$, recent diffusion-based SR approaches (Yang et al., 2023b; Li et al., 2024b; Xie et al., 2024; Du et al., 2023) define an embedding function $\text{Emb}(z)$ that maps discrete items into a continuous space.
+The transition process of $z$ to latent continuous vectors $e_{0}$ is defined as:
 (5) $q_{\phi}(e_{0}|z)=\mathcal{N}(e_{0};\text{Emb}(z),\mathbf{0}),$
-where $\phi$ represents the learnable parameters in $\text{Emb}(z)$. As for the reverse process, we define the predicted distribution of $z$ as:
+where $\phi$ represents the learnable parameters in $\text{Emb}(z)$. In this paper, the corresponding reverse process, which maps vectors in the embedding space back to items is defined as follows:
 (6) $p_{\phi}(z|e_{0})=\arg\max\limits_{z}\text{cos}(e_{0},\text{Emb}(\mathbb{Z})),$
 where $\text{cos}(\cdot)$ is the cosine similarity between $e_{0}$ and each item embedding of $\text{Emb}(\mathbb{Z})$.
-However, the transition distribution $p_{\phi}(z|e_{0})$ lacks a direct analytical formula.
-After defining $p_{\phi}$ and $q_{\phi}$, we extend the ELBO in Equation 7 to include discrete item $z$ as:
-(13) $\arg\max\limits_{\theta}\log p_{\theta}(z)$ $\approx\arg\min\limits_{\phi,\theta}\mathbb{E}_{q_{\phi}(e_{0:T}|z)}\left[\log\frac{q_{\phi}(e_{0:T}|z)}{p_{\phi,\theta}(z,e_{0:T})}\right],$
-and the training objective in Equation 9 is rewritten as:
+After defining $p_{\phi}$ and $q_{\phi}$, the training objective in Equation 3 can be extended to including $z$ as:
 (7) $\mathcal{L}_{\text{elbo}}(z)=\mathbb{E}_{q_{\phi}(e_{0:T}|z)}\left(\mathcal{L}_{T}+\mathcal{L}_{t}+\mathcal{L}_{z}\right),\rm where\$
 […]
-There are three terms of objectives in $\mathcal{L}_{\text{elbo}}$: the prior matching term $\mathcal{L}_{T}$, the denoising term $\mathcal{L}_{t}$ and the domain transition term $\mathcal{L}_{z}$.
-The prior matching term $\mathcal{L}_{T}$ is similar to the same term in Equation 9, which requires the forward process to transit the $e_{0}$ into Gaussian noise $p_{\theta}(e_{T})$.
-Continuous diffusion models (Luo, 2022; Ho et al., 2020; Ho and Salimans, 2022a; Kong et al., 2021) omit this term since there are no associated trainable parameters.
-This term is essential in $\mathcal{L}_{\text{elbo}}$ due to the correlation among $z$, $e_{0}$ and $\phi$.
-This term is equivalent to minimizing $\|e_{0}\|^{2}$.
-As the second term $\mathcal{L}_{t}$ require $f_{\theta}(e_{t},t)$ and $e_{0}$ as close as possible, we transfer the objective $\|e_{0}\|^{2}$ to $\|\sum_{t=1}^{T}f_{\theta}(e_{t},t)\|^{2}$ for fully optimize the DNN models.
-The embedding entropy term $\mathcal{L}_{z}$ can be regarded as a KL-divergence $D_{\text{KL}}(q_{\phi}(e_{0}|z)||p_{\phi}(z|e_{0}))$.
-Although $q_{\phi}(e_{0}|z)$ is a linear transition, the distribution of $p_{\phi}(z|e_{0})$ is unknown.
-Therefore, we replace $D_{\text{KL}}(q_{\phi}(e_{0}|z)||p_{\phi}(z|e_{0}))$ by $D_{\text{KL}}(q(z)||p_{\phi}(z))=D_{\text{KL}}(q(z)||p_{\phi}(z))$.
-The simplified objective of $\mathcal{L}_{\text{elbo}}$ is:
+There are three terms of objectives: the prior matching term $\mathcal{L}_{T}$, the denoising term $\mathcal{L}_{t}$ and the domain transition term $\mathcal{L}_{z}$.
+Different diffusion-based works may have different approaches to optimize $\mathcal{L}_{\text{elbo}}$, while in this paper, our optimization is based on (Li et al., 2022; Gong et al., 2023).
+Specifically, The $\mathcal{L}_{T}$ is similar to the same term in Equation 3, which requires the forward process to transit the $e_{0}$ into Gaussian noise $p_{\theta}(e_{T})$.
+While vanilla diffusion models (Luo, 2022; Ho et al., 2020; Ho and Salimans, 2022a; Kong et al., 2021) omit this term since there are no associated trainable parameters, this term is essential in Equation 7, since $e_{0}$ are calculated from $z$ and $\phi$.
+Following (Li et al., 2022), this term can be simplified as $\|e_{0}\|^{2}$.
+We further transfer the objective $\|e_{0}\|^{2}$ to $\sum_{t=1}^{T}\|f_{\theta}(e_{t},t)\|^{2}$ since the second term $\mathcal{L}_{t}$ requires $f_{\theta}(e_{t},t)$ and $e_{0}$ as close as possible.
+As for $\mathcal{L}_{z}$, it can be regarded as a KL-divergence $D_{\text{KL}}(q_{\phi}(e_{0}|z)||p_{\phi}(z|e_{0}))$.
+However, since $p_{\phi}(z|e_{0})$ is a argmax function that lacks a direct analytical formula, we approximate $D_{\text{KL}}(q_{\phi}(e_{0}|z)||p_{\phi}(z|e_{0}))$ by $D_{\text{KL}}(q(z)||p_{\phi}(z))$, where $q(z)$ is the ground truth distribution and $p_{\phi}(z)$ is the generated distribution of item.
+Thus, the simplified objective of $\mathcal{L}_{\text{elbo}}$ is:
 (9) $\mathcal{L}_{\text{elbo}}(z)=\mathbb{E}_{q_{\phi}(e_{0:T}|z)}\left(\mathcal{L}_{T}+\mathcal{L}_{t}+\mathcal{L}_{z}\right),\rm where\$
-(17) $\mathcal{L}_{T}$ $=\|f_{\theta}(e_{t},t)\|^{2},$
-(18) $\mathcal{L}_{t}$ $=\|f_{\theta}(e_{t},t)-e_{0}\|^{2},$
-(19) $\mathcal{L}_{z}$ $=D_{\text{KL}}(q_{\phi}(z)||p_{\phi}(z))$
+(10) $\mathcal{L}_{T}$ $=\textstyle\sum\nolimits_{t=1}^{T}\|f_{\theta}(e_{t},t)\|^{2},$
+(11) $\mathcal{L}_{t}$ $=\textstyle\sum\nolimits_{t=1}^{T}\|f_{\theta}(e_{t},t)-e_{0}\|^{2},$
+(12) $\mathcal{L}_{z}$ $=D_{\text{KL}}(q_{\phi}(z)||p_{\phi}(z)),$
+3. Dual Conditional Diffusion Model
+In this section, we present a comprehensive introduction of our Dual Conditional Diffusion Model (DCRec) as shown in Figure 2.
+First, we introduce the theoretical derivation of the implicit conditional forward process and the explicit conditional reverse process.
+Then, to bridge the gap between diffusion theory and a practical model, we introduce the Dual Conditional Diffusion Transformer (DCDT), our denoising network $f_{\theta}$ for diffusion SR is shown in Figure 3.
+Finally, we explore efficient inference for DCRec.
+3.1. Implicit Conditional Forward Process
+Implicit conditioning provides indirect guidance for the denoising model by integrating the user’s historical information into the noisy input data.
+Due to the input being calculated by Equation 2, this conditioning effects the model in the diffusion forward process.
+Previous implicit conditional methods tend to encode the user’s historical behavior sequences into compact feature representations, which are then used to guide the reverse diffusion process (Yang et al., 2023b; Li et al., 2024a).
+However, compressing the historical behavior sequences into a single vector might lead to the loss of important sequential information due to the sparse nature of SRSs (He and McAuley, 2016b).
+In our approach, we avoid this compression by concatenating the uncompressed historical behavior sequences with the target item and adding noise to the sequence.
+We believe that this allows the diffusion model to generate a more robust representation of the target item, as it can better capture detailed sequential patterns and contextual information from the user’s history.
+Here, we use $S_{0}$ to denote the concatenated sequence of the history item sequence $H_{0}$ and the target $e_{0}$, where $S_{0}=\text{Concat}(H_{0},e_{0})$.
+The input of the denoising network $f_{\theta}$ is expanded from $e_{0}$ to $S_{0}$.
+Thus, we rewrite Equation 2 as follows:
+(13) $q(S_{t}|S_{t-1})$ $=\mathcal{N}(S_{t};\sqrt{\alpha_{t}}S_{t-1},(1-\alpha_{t})\mathbf{I}),$
+Meanwhile, the denoising DNN $f_{\theta}$ in the reverse process is extended to support sequential input as follows:
+(14) $\bar{S}_{0}=f_{\theta}(S_{t},t),$
+where $\bar{S}_{0}$ is the recovered target item.
+Figure 3.
+The design of DCDT.
+The black box refers to the main architecture, the red box refers to the details within the transformer block, and the yellow box refers to the details of the CondLN module.
+3.2. Explicit Conditional Reverse Process
+Explicit conditioning provides direct guidance for the denoising model by influencing the reverse denoising process in Equation 1, thereby recovering the target item underlying the user’s preferences.
+Although this conditioning has proven promising in image synthesis tasks (Rombach et al., 2022; Podell et al., 2024), most recent diffusion-based SR methods (Yang et al., 2023b; Li et al., 2024b) primarily focus on implicit conditioning, without fully exploring the potential of explicit conditioning.
+Unlike implicit conditioning which may lose part of the sequential information in the user’s behavior sequence, explicit conditioning directly introduces the entire sequence of historical item embeddings.
+After expanding the concatenated input $S_{t}$ in Section 3.1, we rewrite standard reverse transition Equation 1 with item preference $H_{0}$ as an explicit conditional signal for the denoising process:
+(15) $p_{\theta}(S_{t-1}|S_{t},H_{0})=\mathcal{N}(S_{t-1};\mu_{\theta}(S_{t},t,H_{0}),\Sigma_{\theta}(S_{t},t,H_{0})),$
+Combining implicit and explicit conditioning, the denoising DNN $f_{\theta}$ in Equation 14 can be further rewritten as:
+(16) $\bar{S}_{0}=f_{\theta}(S_{t},t,H_{0}),$
+After deriving the forward and reverse process as Equations 13 and 15, implementing a practical model remains challenging.
+In the next section, we introduce a novel network architecture designed to effectively integrate both types of conditioning.
 3.3. Dual Conditional Diffusion Transformer
-In this section, we provide a detailed explanation of the Dual Conditional Diffusion Transformer (DCDT) in DCRec, which is a robust denoising transformer built within the framework in Section 4.1.
-Previous methods in diffusion-based SR have primarily employed either implicit or explicit conditioning mechanisms to incorporate user behavioral data into the recommendation process (Yang et al., 2023b; Li et al., 2024a; Xie et al., 2024).
-Almost implicit conditional approaches compress the user’s historical behavior sequence into a compact representation vector, which is then used to guide the reverse diffusion process (Lin et al., 2024b; Li et al., 2024a).
+We introduce the Dual Conditioned Diffusion Transformer (DCDT) in DCRec, a robust denoising network incorporating both implicit and explicit conditioning.
+Previous implicit conditional approaches (Yang et al., 2023b; Li et al., 2024a; Xie et al., 2024) compress the user’s historical behavior sequence into a compact representation vector, which is then used to guide the reverse diffusion process (Lin et al., 2024b; Li et al., 2024a).
 While this method efficiently captures global user preferences, it may inadvertently oversimplify complex user behaviors, losing important sequential and temporal dynamics inherent in the user’s interactions.
 This compression can lead to a loss of valuable information about the order and context of past actions, which are crucial for accurate SRs.
-Explicit conditioning methods (Lin et al., 2024a), on the other hand, skip the feature compression step and directly introduce the entire sequence of historical item embeddings during the reverse generation process.
+Explicit conditioning methods, on the other hand, skip the feature compression step and directly introduce the entire sequence of historical item embeddings during the reverse generation process.
 This allows the model to access detailed step-by-step interactions and capture fine-grained sequential dependencies.
-However, explicit conditioning on the whole historical information, these methods can be more sensitive to noise and irrelevant actions within the user’s behavior sequence, which may hinder the recommendation quality (He and McAuley, 2016a, b).
-To tackle these limitations, we propose the Dual Conditioned Diffusion Transformer (DCDT) as shown in Figure 3, a novel approach that integrates both implicit and explicit conditioning strategies.
-By combining the representation of the user’s global preferences with the direct utilization of detailed historical behaviors, DCDT  leverages the strengths of both methods.
-The implicit component captures global user interests and long-term behavioral patterns, providing a stable guide for the diffusion process.
-Simultaneously, the explicit component preserves sequential complexities and temporal dynamics by referencing specific past interactions during generation.
-DCDT  employs a dual conditional diffusion process that incorporates an implicit conditional historical forward process and an explicit conditional reverse process, leading to more accurate recommendations.
-As shown in Figure 2(b), the dual conditional trajectory (green line) in the reverse diffusion process merges these two guiding signals, preventing the denoising trajectory from being overly influenced by noise or irrelevant historical items.
-This effectively narrows the trajectory toward the correct target item, improving the precision of the recommendation.
-4.2.1. Implicit Conditional Historical Forward Process:
-Previous implicit conditional methods tend to encode the user’s historical behavior sequences into compact feature representations, which are then used to guide the reverse diffusion process (Yang et al., 2023b; Li et al., 2024a).
-However, compressing the historical behavior sequences into a single vector might lead to the loss of important sequential information due to the sparse nature of SRSs (He and McAuley, 2016b).
-In our approach, we avoid this compression step.
-Instead, we combine the uncompressed historical behavior sequences with the target item and add noise to this combined input before feeding it into the diffusion model.
-We believe that this allows the diffusion model to generate a more robust representation of the target item, as it can better capture detailed sequential patterns and contextual information from the user’s history.
-For convenience, for our implicit conditional diffusion models, we use $S_{0}$ to donate the concatenated sequence of the history item sequence $H_{0}$ and the target $e_{0}$, where $S_{0}=\text{Concat}(H_{0},e_{0})$.
-Thus, we continuously incorporate Gaussian noises into the whole item sequences $S_{0}$ rather than the target item embedding $e_{0}$ as follows:
-(20) $q(S_{t}|S_{t}-1)$ $=\mathcal{N}(S_{t};\sqrt{\alpha_{t}}S_{t-1},(1-\alpha_{t})\mathbf{I})$
-(21) $=\sqrt{\alpha_{t}}S_{0}+\sqrt{1-\alpha_{t}}S_{t-1}$
-Based on the superposition of Gaussian distribution and Equation 21, we can calculate the transition from $S_{0}$ to $S_{t}$ as:
-(22) $q(S_{t}|S_{0})=\sqrt{\bar{\alpha}_{t}}S_{0}+\sqrt{1-\bar{\alpha}_{t}}S_{T},\ S_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
-where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, and $S_{t}$ denotes the item embeddings in the forward step $t$.
-We further add explicit conditioning by adding $H_{0}$ with a scaling factor $\delta$ to $S_{T}$ and rewrite Equation 22 as:
-(23) $q(S_{t}|S_{0})=\sqrt{\bar{\alpha}_{t}}S_{0}+\sqrt{1-\bar{\alpha}_{t}}(S_{T}+\delta H_{0}),$
-Therefore, in the ELBO-based training objective, we can rewrite Equation 17 and Equation 18 as:
-[…]
-Compared to existing implicit diffusion models-based SR methods (Yang et al., 2023b), our method adds noising the entire user sequence rather than just the target item.
-This approach allows the following reverse process to explicitly leverage the user’s interaction history when generating the target item that aligns with the user’s interests, leading to more accurate and relevant recommendations.
-Figure 3. Our DCDT  comprises explicit conditional layer normalization, multi-head self-attention, and an explicit cross-attention.
-4.2.2. Explicit Conditional Reverse Process:
-In the context of SR, the restoration process reverses the transition modeling by gradually eliminating the noise and recovering the original sequential patterns of user behavior, thereby restoring the data distribution underlying the user’s item preferences.
-Meanwhile, cross attention as an explicit conditional method has proved effective in classic conditional generation tasks (Rombach et al., 2022; Podell et al., 2024).
-Therefore, we further introduce the user’s item preferences $H_{0}$ as an explicit conditional signal to the denoising transformer $f_{\theta}$ in Equation 24 as:
-(25) $\hat{S}_{0}=f_{\theta}(S_{t},t,H_{0})$
-Firstly, we introduce the Conditional Layer Normalization, abbreviated as CondLN.
+However, explicit conditioning on the whole historical information can be more sensitive to noise and irrelevant actions within the user’s behavior sequence, which may hinder recommendation quality (He and McAuley, 2016a, b).
+To address these limitations, we jointly leverage both implicit and explicit conditioning.
+We propose DCDT.
+Figure 3 shows an overview of the complete DCDTarchitecture.
+In the following, we describe the architecture design and the loss function of DCDT.
+3.3.1. Input Concatenation
+The main input to DCDT  is a sequential noisy embedding $S_{t}$, concatenated with noisy user’s history sequence $S_{t}$ and noisy target item $e_{t}$.
+This concatenation provides implicit guidance for recover the pure target item.
+Following the concatenation, we apply a standard positional embedding as other transformer-based SRs (Kang and McAuley, 2018; Hidasi and Karatzoglou, 2018; Yuan et al., 2019; Sun et al., 2019; Zhou et al., 2020a).
+3.3.2. Conditional Layer Normalization (CondLN):
 In our SR setting, we hope to extend traditional layer normalization to our explicit conditional denoising process.
 Therefore, we replace the element-wise learnable parameters gain $g$ and bias $b$ of affine transformation in the original Layer Normalization (Lei Ba et al., 2016) by integrating user interaction sequence $S_{0}$ and diffusion step embedding $t$.
-The CondLN can be formulated as:
-(26) $\bar{S}_{t}=g^{\prime}(t,H_{0})\times\frac{S_{t}-\mathbb{E}(S_{t})}{\sqrt{\text{Var}(S_{t})+\epsilon}}+b^{\prime}(t,H_{0})$
-where $\mathbb{E}(\cdot)$ and $\text{Var}(\cdot)$ denote the mean and standard variance of $S_{t}$, and $\epsilon$ denotes the small constant for numerical stability.
-$g^{\prime}(\cdot)$, $b^{\prime}(\cdot)$ are the MLP function conditioned by $H_{0}$ and $t$.
+CondLN can be formulated as:
+(17) $\bar{S}_{t}=g^{\prime}(t,H_{0})\times\frac{S_{t}-\mathbb{E}(S_{t})}{\sqrt{\text{Var}(S_{t})+\epsilon}}+b^{\prime}(t,H_{0}),$
+where $\mathbb{E}(\cdot)$ and $\text{Var}(\cdot)$ denote the mean and variance of $S_{t}$, and $\epsilon$ denotes a small constant for numerical stability.
+$g^{\prime}(\cdot)$, $b^{\prime}(\cdot)$ are MLP functions conditioned by $H_{0}$ and $t$.
 We scale the normalized $\text{Var}(S_{t})$ and shift it with $\mathbb{E}(S_{t})$.
-In this way, CondLN explicitly affects the generation process with user interaction history $H_{0}$.
-This allows for more adaptive and context-aware normalization compared to standard layer normalization, which is crucial for capturing the intricate dependencies in SR tasks.
-Next, we input this sequence embedding $\bar{S}_{t}$ into a self-attention module (SA):
-(27) $\mathbf{\bar{S}_{t}}=\text{SA}(\bar{S}_{t})=\text{Softmax}\left(\frac{(\bar{S}_{t}\mathbf{W}^{Q})(\bar{S}_{t}\mathbf{W}^{K})^{\top}}{\sqrt{d}}\right)(\bar{S}_{t}\mathbf{W}^{V})\in\mathbb{R}^{(n-1)\times d}$
+In this way, CondLN explicitly affects the generation process with the user interactions $H_{0}$.
+This allows for more adaptive and context-aware normalization compared to standard layer normalization, which is crucial for capturing the intricate dependencies.
+3.3.3. Implicit Self-Attention:
+Following the CondLN, we input the sequence embedding $\bar{S}_{t}$ into a self-attention module (SA):
+(18) $\bar{S}_{t}=\text{SA}(\bar{S}_{t})=\text{Softmax}\left(\frac{(\bar{S}_{t}\mathbf{W}^{Q})(\bar{S}_{t}\mathbf{W}^{K})^{\top}}{\sqrt{d}}\right)(\bar{S}_{t}\mathbf{W}^{V})\in\mathbb{R}^{(n-1)\times d},$
 where $\mathbf{W}^{Q}$,$\mathbf{W}^{K}$,$\mathbf{W}^{V}$ are learnable weight parameters, $d$ denotes the dimension of embedding.
-After obtaining the self-attention output, then we leverage the Cross-attention layer (CA):
-(28) $\bar{S}_{t}=\text{CA}(\bar{S}_{t},\mathbf{H_{0}})=\text{Softmax}\left(\frac{(\bar{S}_{t}\mathbf{W}^{Q})(\mathbf{H_{0}}\mathbf{W}^{K})^{\top}}{\sqrt{d}}\right)(\mathbf{H_{0}}\mathbf{W}^{V})$
+The self-attention module captures sequential features from the noisy input $\bar{S}_{t}$ and extracts the high-level representations implicit in the user preference.
+Since there is no extra conditioning input, this module is indirectly guided by the input sequence.
+3.3.4. Explicit Cross-Attention:
+After obtaining the self-attention output and applying layer normalization, we utilize the cross-attention module (CA) with the pure historical embedding $H_{0}$:
+(19) $\bar{S}_{t}=\text{CA}(\bar{S}_{t},\mathbf{H_{0}})=\text{Softmax}\left(\frac{(\bar{S}_{t}\mathbf{W}^{Q})(\mathbf{H_{0}}\mathbf{W}^{K})^{\top}}{\sqrt{d}}\right)(\mathbf{H_{0}}\mathbf{W}^{V}),$
 where $\mathbf{W}^{Q}$, $\mathbf{W}^{K}$, $\mathbf{W}^{V}$ are learnable parameters.
-Due to the sparse interaction in SR (He and McAuley, 2016b), we adapt $\bar{S}_{t}$ as query and the $H_{0}$ (pure historical embeddings) as key, value.
-$H_{0}$ can provide more rich user interaction information than ambiguous $H_{t}$.
-Hence, we select a cross-attention architecture as the denoising decoder instantiation capable of taking the entire sequence representation and corresponding step indicator $t$ as input.
+Due to the sparse interactions in SR (He and McAuley, 2016b), we adapt $\bar{S}_{t}$ as a query and $H_{0}$ as both key and value.
+$H_{0}$ can explicitly provide more rich user interaction information than ambiguous $H_{t}$.
 Hence, we adopt a cross-attention architecture for the denoising decoder, which effectively processes the entire sequence representation along with the corresponding diffusion step indicator $t$ as input.
-Then we endow the traditional layer normalization and feed-forward network to obtain the final user sequence embedding $\hat{S}_{0}$:
-(29) $\bar{S}_{t}=\text{LayerNorm}(\bar{S}_{t}+\text{Dropout}(\text{sublayer}(\bar{S}_{t})))$
-(30) $\hat{S}_{0}=\text{FFN}(\bar{S}_{t})=\text{RELU}(\bar{S}_{t}\times\mathbf{W}+\mathbf{b})$
-After obtaining the $\hat{S}_{0}$, we extract the last item embedding $e_{0}$ of $\hat{S}_{0}$ as the generative target item embedding.
-Then we leverage the variant of Equation 12, to transit continuous $\hat{S}_{0}$ into the discrete item space $\mathbb{Z}$ as $z$, taking the item index $z$ corresponding to the maximal as the final recommendation list to the specific user.
-Finally, the optimization object of the dual conditional diffusion process can be reformulated as:
+3.3.5. Traditional Feed-forward:
+Then we employ another layer normalization and the traditional feed-forward module to obtain the final sequence embedding of the DCDT  block:
+(20) $\bar{S}_{t}=\text{FFN}(\bar{S}_{t})=\text{RELU}(\bar{S}_{t}\times\mathbf{W}+\mathbf{b}),$
+where $\mathbf{W}$ and $\mathbf{b}$ are learnable parameters.
+3.3.6. Network Output and Loss Function:
+After the repeated DCDT  blocks and the final CondLN module, we obtain the predicted target $\bar{S}_{0}=\text{Concat}(\bar{H}_{0},\bar{e}_{0})$.
+In training, we extract the last item embedding $\bar{e}_{0}$ of $\bar{S}_{0}$ as the generative target item embedding.
+Then we leverage the variant of Equation 6, to transit continuous $\bar{S}_{0}$ into the discrete item space $\mathbb{Z}$ as $\bar{z}$, taking the item index $\bar{z}$ corresponding to the maximal probability as the final recommendation list to the specific user.
+Finally, the standard optimization object Equation 9 can be reformulated for our dual conditional diffusion as:
 (21) $\mathcal{L}_{\text{elbo}}(z)=\mathbb{E}_{q_{\phi}(e_{0:T}|z)}\left[\mathcal{L}_{T}+(1-\lambda)\mathcal{L}_{t}+\lambda\mathcal{L}_{z}\right],\rm where\$
-(32) $\mathcal{L}_{T}$ $=\|f_{\theta}(e_{t},t)\|^{2},$
-(33) $\mathcal{L}_{t}$ $=\|f_{\theta}(e_{t},t)-e_{0}\|^{2},$
-(34) $\mathcal{L}_{z}$ $=D_{\text{KL}}(q_{\phi}(z)||p_{\phi}(z))$
-where $\lambda$ donates the balance factor mediating the focus between $\mathcal{L}_{t}$ and $\mathcal{L}_{z}$. Algorithm 1 in Appendix A provides a detailed explanation of DCRec’s training process.
-4.3. Model Inference and Inference Acceleration
-In the Section 4.1 and Section 4.2, we have introduced the optimization process in DCRec  and the training of DCDT.
-The denoising model DCDT  is trained to provide predicted $\hat{S}_{0}$ to approximate $S_{0}$.
-Similar to Equation 2, the inference Markov transition of DCRec  can be formulated as:
-(35) $p_{\theta}(S_{t-1}|S_{t})=\mathcal{N}(S_{t-1};\mu_{\theta}(\hat{S}_{0}(S_{t}),t),\Sigma_{\theta}(\hat{S}_{0}(S_{t}),t))$
-This can be described as a repeated approximation-and-adjust process since each step aims to predict $\hat{S}_{0}(S_{t})$ and provide $S_{t-1}$, which is less noise than $S_{t}$, for the next step.
-The time cost of inference process linearly increases associated with the increase of the total number of steps.
-A natural idea for acceleration is that, if the approximation at the early is accurate enough with no need for adjustment, DCRec  can directly skip several intermediate steps.
+(22) $\mathcal{L}_{T}$ $=\|\bar{S}_{0}\|^{2},$
+(23) $\mathcal{L}_{t}$ $=\|\bar{e}_{0}-e_{0}\|^{2},$
+(24) $\mathcal{L}_{z}$ $=D_{\text{KL}}(q_{\phi}(z)||p_{\phi}(\bar{z})),$
+where $\lambda$ donates the balance factor between $\mathcal{L}_{t}$ and $\mathcal{L}_{z}$. Algorithm 1 provides details of the training process in DCRec.
+Algorithm 1 DCRec  Training
+1:User interacted item embedding set $\{H_{0}\}$, item pool $\mathbb{Z}$, number of diffusion steps $T$, denoising network $f_{\theta}$, balance factor $\lambda$
+2:repeat
+3:     Sample ${H_{0}}$ from $\{H_{0}\}$ and corresponding $z$ from $\mathbb{Z}$
+4:     $S_{0}$leftarrow Concatenate $H_{0}$ and $z$
+5:     Sample $t\sim\mathcal{U}(1,T)$
+6:     $S_{t}$leftarrow Adding noise to $S_{0}$ according to Equation 2
+7:     Obtain predicted $\bar{S}_{0}\leftarrow f_{\theta}(S_{t},t,H_{0})$
+8:     Extract predicted $\bar{e}_{0}$ from $\bar{S}_{0}$
+9:     Obtain predicted $\bar{z}\leftarrow p_{\phi}(\bar{z}|\bar{e}_{0})$
+10:     $\mathcal{L}_{T}$leftarrow $\|\bar{S}_{0}\|^{2}$
+11:     $\mathcal{L}_{t}$leftarrow $\|\bar{e}_{0}-e_{0}\|^{2}$
+12:     $\mathcal{L}_{z}$leftarrow $D_{\text{KL}}(q_{\phi}(z)||p_{\phi}(\bar{z}))$
+13:     Take gradient descent step on$\nabla\left(\mathcal{L}_{T}+(1-\lambda)\mathcal{L}_{t}+\lambda\mathcal{L}_{z}\right)$ to optimize $\theta$ and $\phi$
+14:until converged
+15:optimized $\theta$, $\phi$.
+3.4. Model Inference and Acceleration
+In the Section 3.3, DCDT  is trained to provide predicted $\bar{S}_{0}$ to approximate $S_{0}$.
+During model inference, DCDT  iteratively generates $S_{0}$ according to the total steps $T$.
+$\bar{S}_{0}$ will be transfer to $S_{t-1}$ for the next time step, formulated as $S_{t-1}=\bar{S}_{0}(S_{t})$.
+Based on Equation 15, the inference Markov transition of DCRec  can be reformulated as:
+(25) $p_{\theta}(S_{t-1}|S_{t})=\mathcal{N}(S_{t-1};\mu_{\theta}(\bar{S}_{0}(S_{t}),t),\Sigma_{\theta}(\bar{S}_{0}(S_{t}),t))$
+This can be described as a repeated approximation-and-adjustment process since each step aims to generate $\bar{S}_{0}$ base on the output from the last step.
+$\bar{S}_{0}$ becomes increasingly precise as inference progresses.
+The time cost of the inference process increases linearly in the total number of steps.
+A natural idea for acceleration is that if the approximation at the early stage is accurate enough with no need for adjustment, DCRec  can directly skip several intermediate steps according to the high similarity of $\bar{S}_{0}$.
 Here, we define the number of skipping steps as k.
 We rewrite Equation 25 as:
-(36) $p_{\theta}(S_{t-k}|S_{t})=\mathcal{N}(S_{t-k};\mu_{\theta}(\hat{S}_{0}(S_{t}),t),\Sigma_{\theta}(\hat{S}_{0}(S_{t}),t))$
+(26) $p_{\theta}(S_{t-k}|S_{t})=\mathcal{N}(S_{t-k};\mu_{\theta}(\bar{S}_{0}(S_{t}),t),\Sigma_{\theta}(\bar{S}_{0}(S_{t}),t)).$
 For our DCRec, we observe that DCDT  is robust enough to fulfill Equation 26 at an early stage, which significantly reduces time overhead.
 We will discuss the details in Section 4.5.
-For the pseudo-code of the inference, please refer to Algorithm.
-2 in the Appendix.
-5. Experiment
+Algorithm 2 shows the pseudo-code of the inference procedure.
+Algorithm 2 DCRec  Inference
+1:User interacted item embedding $H_{0}$, implicit conditioning scale factor $\delta$, number of diffusion steps $T$, number of accelerating steps $T^{\prime}$, denoising network $f_{\theta}$
+2:Uniformly sample $T^{\prime}$ steps from $1$ to $T$
+3:Sample $e_{T}~{}\sim\mathcal{N}(\bm{0},\bm{I})$
+4:Sample $H_{T}~{}\sim\mathcal{N}(\bm{0},\bm{I})$
+5:Implicit conditioning $H_{T}\leftarrow H_{T}+\delta H_{0}$
+6:$S_{T}\leftarrow$ Concatenate $H_{T}$ and $e_{T}$
+7:for $t\leftarrow T^{\prime}\dots 1$  do
+8:     Obtain predicted ${S}_{0}\leftarrow f_{\theta}(S_{t},t,H_{0})$
+9:     Transit to $S_{t-1}\leftarrow q(S_{t-1}|{S}_{0})$
+10:Extract $e_{0}$ from $S_{0}$
+11:obtain $z\leftarrow p_{\phi}(z|e_{0})$
+12:$z$
+4. Experiments
 In this section, we conduct a comprehensive experimental study to address the following research questions (RQs):
-$\bullet$ RQ1: How does the performance of DCRec  compare with other baselines across the datasets in different experiment settings?
-$\bullet$ RQ2: What is the impact of different components (e.g., reconstruction loss, ranking loss, regularization loss, and DCDT  modules) within the DCRec  on overall performance?
-$\bullet$ RQ3: How does DCRec  perform in terms of inference efficiency?
+  • RQ1: How does the performance of DCRec  compare with other baselines across the datasets in different experiment settings?
+  • RQ2: What is the impact of different components (e.g., concatenated input, reconstruction loss, ranking loss, regularization loss, and DCDT  modules) within the DCRec  on overall performance?
+  • RQ3: What is the different impact of implicit, and explicit conditioning in DCRec?
+  • RQ4: How does DCRec  perform in terms of inference efficiency?
+  • RQ5: How does the noise concatenated input effect the diffusion model performance?
 4.1. Experimental Settings
-$\bullet$ Datasets.
-We evaluate our proposed DCRec  on three publicly available datasets: Amazon 5-core Beauty, Amazon 5-core Toys, and Yelp, following the experimental setup of previous works (Zhou et al., 2020a; Chen et al., 2022a; Li et al., 2024b).
-Detailed descriptions of these datasets can be found in Appendix B due to space constraints.
-$\bullet$ Baselines.
-To evaluate the performance of our model, we select various baselines covering three types of mainstream methods in SR tasks, e.g., representative SR methods, contrastive learning SR methods, and diffusion-based SR methods.
-The detailed information is listed in the Appendix C.
-$\bullet$ Implementation Details and Parameter Settings.
-To ensure fair comparisons, all the baselines and ours are implemented based on the popular recommendation framework RecBole (Zhao et al., 2021) under the same settings and evaluated consistently.
-The detailed information is listed in the Appendix F.
-$\bullet$ Evaluation Metrics.
-We use two commonly used metrics including Normalized Discounted Cumulative Gain @K (NDCG@K), Hit Ratio @K (HR@K) to evaluate the performance of all compared methods.
-For all the baseline models, we generate the ranking list of items for each testing interaction.
-Both of them are applied with $K$ chosen from {5, 10}, evaluating the ranking results using all items in the dataset as candidates.
-Meanwhile, we follow the setting in (Wang et al., 2019b), using the paired t-test with p¡0.05 for the significance test.
-5.2. Overall Performance (RQ1)
+4.1.1. Datasets.
+We evaluate our proposed DCRec  on three publicly available datasets: Amazon 5-core Beauty, Amazon 5-core Toys, and Yelp, following the experimental setup of previous works (Zhou et al., 2020b; Chen et al., 2022a; Li et al., 2024b).
+We evaluate our proposed DCRec  on three publicly accessible datasets in different experiment settings.
+Table 1 provides a summary of the statistics for each dataset.
+  • Amazon Beauty & Amazon Toys: We choose three representative sub-datasets from Amazon datasets: Beauty and Toys and keep the ‘5-core’ datasets (Zhou et al., 2020b) (Rendle et al., 2010)(Chen et al., 2022a), which filter out user-item interaction sequences with length less than 5.
+  • Yelp: Yelp is a well-known business recommendation dataset. Following (Xie et al., 2022; Chen et al., 2022b), we only retain the transaction records after Jan. 1st, 2019 in our experiment.
+For reproducibility, we follow the commonly used benchmark setting of RecBole (Zhao et al., 2021) to set up our experiments.
+For each user, we first discard duplicated interaction sequences and sort the items in each user’s sequence chronologically by their timestamp.
+The maximum length of interaction sequences is set to 50.
+If there are more than 50 interactions in a sequence, we adopt the latest 50 interactions.
+If the number of interactions in a sequence is less than 50, we make it up to 50 by padding virtual items with the ID of 0.
+Following the common practice in SR, we leave the interactions happening at the latest time as the test dataset and the interactions at the second latest time as the validation dataset (Chen et al., 2022a; Zhou et al., 2020c).
+Table 1. Statistics of the datasets after preprocessing.
+Dataset Beauty Toys Yelp
+# Users 22,363 19,412 30,499
+# Items 12,101 11,924 20,068
+# Avg. Actions / User 8.9 8.6 10.4
+# Avg. Actions / Item 16.4 14.1 15.8
+# Actions 198,502 167,597 317,182
+Sparsity 99.93% 99.93% 99.95%
 Table 2.
@@
 Dataset Metric GRU4Rec Caser SARSRec BERT4Rec S3Rec CL4SRec ICLRec DreamRec DiffuRec SdifRec DCRec Improv.
-Beauty HR@5 0.0363 0.0374 0.0337 0.0209 0.0189 0.0423 0.0493 0.0498 0.0559 0.0609 $\textbf{0.0630}^{*}$ 3.45%
+Beauty HR@5 0.0363 0.0251 0.0374 0.0209 0.0189 0.0423 0.0493 0.0498 0.0559 0.0609 $\textbf{0.0630}^{*}$ 3.45%
 HR@10 0.0571 0.0601 0.0635 0.0393 0.0307 0.0694 0.0737 0.0698 0.0783 0.0819 $\textbf{0.0859}^{*}$ 4.88%
@@
 NDCG@10 0.0124 0.0113 0.0136 0.0167 0.0092 0.0197 0.0217 0.0152 0.0313 0.0201 $\textbf{0.0338}^{*}$ 7.99%
+4.1.2. Baselines.
+To evaluate the performance of our model, we select various baselines covering three types of mainstream methods in SR tasks, e.g., representative SR methods, contrastive learning SR methods, and diffusion-based SR methods.
+To evaluate the performance of our model DCRec, we select various representative and/or state-of-the-art recommendation models as baselines.
+  • GRU4Rec (Hidasi and Karatzoglou, 2018): This model uses the GRU to model the user interaction sequence and gives the final recommendation.
+  • Caser (Yuan et al., 2019): CNN-based approach for SR.
+  • SARSRec (Kang and McAuley, 2018): This model employs a single-direction Transformer with a masked encoder to capture explicit correlations between items.
+  • Bert4Rec (Sun et al., 2019): replaces the next-item prediction with a Cloze task to fuse information between an item (a view) in a user behavior sequence and its contextual information.
+  • S3Rec (Zhou et al., 2020a): uses SSL to capture correlation-ship among items, sub-sequences, and associated attributes from the given user behavior sequence.
+Its modules for mining on attributes are removed because we don’t have attributes for items.
+  • CL4SRec (Xie et al., 2022): fuses contrastive SSL with a Transformer-based SR model.
+  • ICLRec (Chen et al., 2022b): leverages latent user intents, learned from unlabeled behavior sequences, to optimize SR (SR) models using contrastive self-supervised learning.
+  • DreamRec (Yang et al., 2023b): uses the historical interaction sequence as conditional guiding information for the diffusion model to enable personalized recommendations.
+  • SdifRec (Xie et al., 2024): introduces the Schrödinger Bridge into diffusion-based SR, replacing the Gaussian prior with the user’s state.
+  • DiffuRec (Li et al., 2024b): incorporates a diffusion model into the field of SR, using a Transformer backbone to reconstruct target item representations based on the user’s historical interaction behaviors.
+4.1.3. Implementation Details and Parameter Settings.
+To ensure fair comparisons, all the baselines and ours are implemented based on the popular recommendation framework RecBole (Zhao et al., 2021) under the same settings and evaluated consistently.
+We implement our model and baselines using Pytorch (Paszke et al., 2019) and conduct our experiments on an NVIDIA RTX 3090 GPU with 24G memory.
+For a fair comparison, we carefully tuned the model parameters of all the baselines on the validation set.
+Then we choose the model parameters that achieve the best performance on the validation set to compare.
+For our DCRec, based on the hyperparameter analysis, we set the batch size to 512, the item embedding dimension to 128, learning rate of $34\times 1e-$ 3, maximum sequence length to 50, $\beta$ to 0.02, the dropout rate to 0.1, the embedding dropout to 0.4, the hidden factor to 128, $\delta$ to 0.1, the number of transformer blocks to 4, and the number of diffusion steps $T$ to 50.
+For $\lambda$, we adopt $\lambda_{\text{smooth}}$ for Beauty, Toys and Yelp datasets.
+For detailed settings and further analysis of $\lambda$ , please refer to Section 4.3.2.
+Besides, we choose Adam (Kingma and Ba, 2015) as our optimization strategy and use early stopping with a patience of 5 to prevent overfitting.
+4.1.4. Evaluation Metrics.
+We use two commonly used metrics including Normalized Discounted Cumulative Gain @K (NDCG@K), Hit Ratio @K (HR@K) to evaluate the performance of all compared methods.
+For all the baseline models, we generate the ranking list of items for each testing interaction.
+Both of them are applied with $K$ chosen from {5, 10}, evaluating the ranking results using all items in the dataset as candidates.
+Meanwhile, we follow the setting in (Wang et al., 2019b), using the paired t-test with p¡0.05 for the significance test.
+4.2. Overall Performance (RQ1)
 In this section, Table 2 reports the comparison results between our method and 10 different baseline methods on three datasets.
@@
 The results compare four configurations: single attention (Single-attn) with and without Conditional Layer Normalization (CondLN) and dual attention (Dual-attn) with and without CondLN.
-Single-attn uses only a self-attention block where the noised sequence, diffusion step, and history embeddings are directly combined.
-In contrast, Dual-attn integrates both self-attention and cross-attention layers to capture interactions between the noised sequence and user history.
-CondLN conditions the LN process on the noised sequence embedding $S_{t}$, diffusion time embedding $t$ adapting to personalized user behaviors.
+Single-attn uses only a self-attention block where the noisy sequence, diffusion step, and history embeddings are directly combined.
+In contrast, Dual-attn integrates both self-attention and cross-attention layers to capture interactions between the noisy sequence and user history.
+CondLN conditions the LN process on the noisy sequence embedding $S_{t}$, diffusion time embedding $t$ adapting to personalized user behaviors.
 The results show that both CondLN and dual attention effectively boost performance with highest HR@5 performance when combining them.
@@
 As shown in Table 3, DCRec  consistently outperforms both ICDM and ECDM, demonstrating that jointly modeling implicit and explicit historical conditions is more effective than relying on either one individually.
+Meanwhile, compared to implicit conditional diffusion baselines in Table 2, our ICDM has shown competitive performance on Beauty and Yelp, and better performance on Toys.
 Table 3. Performance comparison among Implicit Conditional DMs (ICDM), Explicit Conditional DMs (ECDM), and our DCDT. means the improvement is significant at $p<0.05$.
@@
 4.3.2. Hyper-parameters Analysis.
-For a more nuanced understanding, we select certain sensitive hyper-parameter $\lambda$, and explore three different $\lambda$ settings as shown in Figure 6.
-Figure 6. Figure (a) depicts three types of $\lambda$ and Figure (b) reports the corresponding HR@5 results on Amazon Beauty and Toys datasets. The $\lambda_{smooth}$ reports superior performance.
 The loss balance factor $\lambda(\cdot)$ influences DCRec  performance by mediating the focus between recommendation and reconstruction tasks.
 While an increase in $\lambda$ initially bolsters performance by prioritizing embedding denoising, too high a value risks neglecting the core recommendation task, undermining overall performance.
-$\lambda_{fix}$ denotes experimental constant, $\lambda_{decline}$ denotes handcrafted function, $\lambda_{smooth}$ denotes training epoch related function, respectively.
-We can find that $\lambda_{smooth}$ can achieve the best performance.
-Due to the page limit, the details of $\lambda$ and other hyper-parameters please refer to Appendix E.
-5.4. Model Efficiency Analysis (RQ3)
+For a more nuanced understanding, we explore three different $\lambda$ settings as shown in follows:
+$\lambda_{\text{fix}}&=c,\\ \lambda_{\text{decline}}&=\text{Max}-g_{d}(\text{ep})+\text{Min},\\ \lambda_{\text{smooth}}&=\text{Max}*g_{s}(\text{ep})+\text{Min},\\ \end{aligned}\right.$
+where, $\lambda_{fix}$ denotes an experimental constant $c$.
+$\lambda_{decline}$ and $\lambda_{smooth}$ denotes epoch-related linear function $g_{d}(\text{ep}$ and epoch-related power function $g_{s}(\text{ep})$, respectively, with upper bound Max and lower bound Min.
+We evaluate these three types of $\lambda$ with setting: $c=0.1$, $\{\text{Max},\text{Min}\}\!=\!\{0.2,0.003\}$, $g_{d}\!=\!(0.2-\text{ep}/400)$, and $g_{s}\!=\!(1-\text{ep}/150)^{9}$ as shown in Figure 6(a).
+The results in Figure 6(b) shows that $\lambda_{smooth}$ can achieve better performance.
+Figure 6. (a) depicts three types of $\lambda$ and (b) reports the corresponding HR@5 results on Amazon Beauty and Toys datasets. The $\lambda_{smooth}$ reports superior performance.
+4.4. Model Conditioning Analysis (RQ3)
+We notice that the implicit components capture long-term behavioral patterns well, while the explicit components preserve temporal dynamics.
+To verify our claim, we provide a comparison experiment on the Amazon Beauty dataset.
+First, we divide the dataset into two subsets based on sequence length: one containing sequences with more than 10 interactions and the other with fewer than 10 interactions.
+Next, we train implicit-only, explicit-only, and DCRec models on the original dataset and evaluate their performance on the two subsets.
+The implicit-only and explicit-only models only using corresponding components in DCDT.
+The HR@5 results are shown in Table 4, illustrating that the implicit component captures long-term behavioral patterns well, while the explicit component preserves temporal dynamics.
+Table 4. The effects of different conditioning approaches on different datasets with varying data length (HR@5).
+Dataset Number of Sequence DCRec Implicit Explicit
+Original 22,363 0.0630 0.0539 0.0585
+Short Seq 17,240 0.0591 0.0488 0.0501
+Long Seq 5,123 0.0639 0.0651 0.0635
+4.5. Model Efficiency Analysis (RQ4)
 Figure 7. Effect of inference step $T$ of our DCRec  on Yelp and Beauty datasets ($T$ denotes the diffusion steps in training while $T^{\prime}$ denotes the steps in inference).
 Figure 8. The impact of sample steps on our DCRec  and DiffuRec on HR@5 and time cost (second) on the Beauty dataset.
-In this section, we investigate the impact of inference step $T$ to answer RQ3.
-We vary diffusion steps $T$ from 0 to $T$ in both Beauty and Toys datasets during inference and show the results in Figure 7.
-Moreover, we conducted a time comparison with DiffuRec (Li et al., 2024b), the most representative method in the baseline.
+In this section, we analyze the network complexity of DCDT  and investigate the impact of inference step $T$ to answer RQ4.
+DCRec  has the same overall network complexity $\mathcal{O}(nd^{2}+dn^{2})$ as DreamRec and DiffuRec, where $n$ denotes the sequence length and $d$ is the embedding dimension.
+Besides, we vary diffusion steps from 0 to $T$ in both Beauty and Toys datasets during inference and show the results in Figure 7.
+We observe that the optimal inference steps vary across different datasets, suggesting that the ideal diffusion step for each dataset is influenced by its item count.
+Moreover, we conducted a time comparison with a strong baseline, DiffuRec (Li et al., 2024b).
 To the fair comparison, we set 30 diffusion steps for training our DCRec  aligning to the DiffuRec as shown in Figure 8.
-Firstly, We can find the best inference steps are vary in different datasets.
-This implies that the optimal diffusion step for each dataset is related to their item count.
-Then we observe that our DCDT ($f_{\theta}$) is optimized powerful enough to fulfill Equation 36 at an early $\hat{t}$-th stage, which significantly reduces time overhead.
+We observe that our DCDT ($f_{\theta}$) is optimized powerful enough to fulfill Equation 26 at an early $\bar{t}$-th stage, which significantly reduces time overhead.
 Compared to DiffuRec, our DCRec’s performance is more stable under different inference steps $T$ settings.
-In conclusion, our DCRec  can reach superior model performance than DiffuRec while maintaining lower inference time cost under different inference step settings.
-This is because our improvement enables the dual conditional mechanism, resulting in both accurate and efficient inference.
-6. Conclusion
-In this paper, we proposed the DCRec  to address key limitations in existing diffusion-based sequential recommender systems.
-By introducing a Markov chain to model item transitions and a dual conditional module to integrate contextual information, our approach improves the consistency of item selection and enhances the generation of target item embedding.
-Extensive experiments on public datasets validate the effectiveness of DCRec, demonstrating its superiority over state-of-the-art methods in both recommendation accuracy and computational efficiency.
-Appendix A Algorithm of DCRec  Training and Inference
-Algorithm 1 DCRec  Training
-1:User interacted item embedding set $\{H_{0}\}$, item pool $\mathbb{Z}$, number of diffusion steps $T$, denoising network $f_{\theta}$, balance factor $\lambda$
-2:repeat
-3:     Sample ${H_{0}}$ from $\{H_{0}\}$ and corresponding $z$ from $\mathbb{Z}$
-4:     $S_{0}$leftarrow Concatenate $H_{0}$ and $z$
-5:     Sample $t\sim\mathcal{U}(1,T)$
-6:     $S_{t}$leftarrow Adding noise to $S_{0}$ according to Equation 23
-7:     Obtain predicted $\hat{S}_{0}\leftarrow f_{\theta}(S_{t},t,H_{0})$
-8:     Extract predicted $\hat{e}_{0}$ from $\hat{S}_{0}$
-9:     Obtain predicted $\hat{z}\leftarrow p_{\phi}(\hat{z}|\hat{e}_{0})$
-10:     $\mathcal{L}_{T}$leftarrow $\|\hat{S}_{0}\|^{2}$
-11:     $\mathcal{L}_{t}$leftarrow $\|\hat{e}_{0}-e_{0}\|^{2}$
-12:     $\mathcal{L}_{z}$leftarrow $D_{\text{KL}}(q_{\phi}(z)||p_{\phi}(\hat{z}))$
-13:     Take gradient descent step on$\nabla\left(\mathcal{L}_{T}+(1-\lambda)\mathcal{L}_{t}+\lambda\mathcal{L}_{z}\right)$ to optimize $\theta$ and $\phi$
-14:until converged
-15:optimized $\theta$, $\phi$.
-Algorithm 2 DCRec  Inference
-1:User interacted item embedding $H_{0}$, implicit conditioning scale factor $\delta$, number of diffusion steps $T$, number of accelerating steps $T^{\prime}$, denoising network $f_{\theta}$
-2:Uniformly sample $T^{\prime}$ steps from $1$ to $T$
-3:Sample $e_{T}~{}\sim\mathcal{N}(\bm{0},\bm{I})$
-4:Sample $H_{T}~{}\sim\mathcal{N}(\bm{0},\bm{I})$
-5:Implicit conditioning $H_{T}\leftarrow H_{T}+\delta H_{0}$
-6:$S_{T}\leftarrow$ Concatenate $H_{T}$ and $e_{T}$
-7:for $t\leftarrow T^{\prime}\dots 1$  do
-8:     Obtain predicted ${S}_{0}\leftarrow f_{\theta}(S_{t},t,H_{0})$
-9:     Transit to $S_{t-1}\leftarrow q(S_{t-1}|{S}_{0})$
-10:Extract $e_{0}$ from $S_{0}$
-11:obtain $z\leftarrow p_{\phi}(z|e_{0})$
-12:$z$
-Appendix B Dataset Preparation
-We evaluate our proposed DCRec  on three publicly accessible datasets in different experiment settings. Table 3 provides a summary of the statistics for each dataset.
-  • Amazon Beauty & Amazon Toys: We choose three representative sub-datasets from Amazon datasets: Beauty and Toys and keep the ‘5-core’ datasets (Zhou et al., 2020a) (Rendle et al., 2010)(Chen et al., 2022a), which filter out user-item interaction sequences with length less than 5.
-  • Yelp: Yelp is a well-known business recommendation dataset. Following (Xie et al., 2022; Chen et al., 2022b), we only retain the transaction records after Jan. 1st, 2019 in our experiment.
-For reproducibility, we follow the commonly used benchmark setting of RecBole (Zhao et al., 2021) to set up our experiments.
-Specifically, for each user, we first discard duplicated interaction sequences and sort the items in each user’s sequence chronologically by their timestamp.
-Furthermore, the maximum length of interaction sequences is set to 50.
-If there are more than 50 interactions in a sequence, we adopt the latest 50 interactions.
-If the number of interactions in a sequence is less than 50, we make it up to 50 by padding virtual items with the ID of 0.
-Following the common practice in sequential recommendations, we leave the interactions happening at the latest time as the test dataset and the interactions at the second latest time as the validation dataset (Chen et al., 2022a; Zhou et al., 2020b).
-Table 3. Statistics of the datasets after preprocessing.
-Dataset Beauty Toys Yelp
-# Users 22,363 19,412 30,499
-# Items 12,101 11,924 20,068
-# Avg. Actions / User 8.9 8.6 10.4
-# Avg. Actions / Item 16.4 14.1 15.8
-# Actions 198,502 167,597 317,182
-Sparsity 99.93% 99.93% 99.95%
-Appendix C Baselines for Comparison
-To evaluate the performance of our model DCRec, we select various representative and/or state-of-the-art recommendation models as baselines.
-  • GRU4Rec (Hidasi and Karatzoglou, 2018): This model uses the GRU to model the user interaction sequence and gives the final recommendation.
-  • Caser (Yuan et al., 2019): CNN-based approach for SR.
-  • SARSRec (Kang and McAuley, 2018): This model employs a single-direction Transformer with a masked encoder to capture explicit correlations between items.
-  • Bert4Rec (Sun et al., 2019): replaces the next-item prediction with a Cloze task to fuse information between an item (a view) in a user behavior sequence and its contextual information.
-  • $S^{3}$ec (Zhou et al., 2020c): uses SSL to capture correlation-ship among items, sub-sequences, and associated attributes from the given user behavior sequence.
-Its modules for mining on attributes are removed because we don’t have attributes for items.
-  • CL4SRec (Xie et al., 2022): fuses contrastive SSL with a Transformer-based SR model.
-  • ICLRec (Chen et al., 2022b): leverages latent user intents, learned from unlabeled behavior sequences, to optimize SR (SR) models using contrastive self-supervised learning.
-  • DreamRec (Yang et al., 2023b): uses the historical interaction sequence as conditional guiding information for the diffusion model to enable personalized recommendations.
-  • SdifRec (Xie et al., 2024): introduce the Schrödinger Bridge into diffusion-based SR, by replacing the Gaussian prior with the user’s current state.
-  • DiffuRec (Li et al., 2024b): incorporates a diffusion model into the field of SR, leveraging a Transformer backbone to reconstruct target item representations based on the user’s historical interaction behaviors.
-Appendix D Ablation Study for Partially Noised Input Sequence
-In this paper, we concatenate the history embedding and target item embedding as a sequential input for our denoising model DCDT.
+In conclusion, our DCRec  shows superior model performance than DiffuRec while maintaining lower inference cost under different step settings.
+This is because our improvement enables the dual conditional mechanism, resulting in both accuracy and efficiency.
+4.6. Ablation Study for Noise Adding (RQ5)
+In this paper, we concatenate the history embedding and target item embedding as a sequential input for our denoising network DCDT.
 As shown in Figure 2, we add noise to the whole input sequence.
@@
 The results are shown in Figure 9.
-It seems that partially adding noise to the target item is worse than adding noise to the whole sequence.
-The possible reason is the unbalanced noised sequence length.
+Partially adding noise to the target item proves less effective than adding noise to the whole sequence.
+The possible reason is the unbalanced noisy sequence length.
 DiffuSeq concatenates the question and answer as an input and only adds noise to the answer sequence.
@@
 However, in our SR task, the history sequence is nearly 50×50$\times$50 × longer than the target.
-Therefore, the strategy of partially noised input is unfeasible for DCRec.
-Figure 9. Performance comparison of history conditioning with noise and without noise.
-Appendix E Details of Hyper-parameters Analysis
-E.1. The Balance Factor $\lambda$
-It’s worth to note that the optimization in Equations 31 and 24 is a complex and parameter-sensitive multi-objective task.
-In this section, we detailed analyze the impacts of three types of different $\lambda$ settings, including fix factor $\lambda_{\text{fix}}$, decline factor $\lambda_{\text{decline}}$ and smoothly decline factor $\lambda_{\text{smooth}}$.
-We formulate them as follows:
-$\lambda_{\text{fix}}&=c\\ \lambda_{\text{decline}}&=\text{Max}-g_{d}(\text{ep})+\text{Min}\\ \lambda_{\text{smooth}}&=\text{Max}*g_{s}(\text{ep})+\text{Min}\\ \end{aligned}\right.$
-where $c$ is a constant and ep is the current training epoch.
-Max and Min denote the upper bound and the lower bound of the decline factor $\lambda_{\text{decline}}$ and $\lambda_{\text{smooth}}$.
-$g_{d}(\text{ep})$ is a decline linear function associated with the current training epoch, while $g_{s}(\text{ep})$ is another decline exponential function also associated with the epoch.
-In Table 4 and Figure 10, we illustrate several example settings of $\lambda$ in our experiments and corresponding results are shown in Table 5.
-We note that those declining settings dramatically improve the model performance compared to the fixed ones.
-The $\lambda_{\text{smooth}}$ settings are feasible for Toys and Beauty datasets, while $\lambda_{\text{decline}}$ settings are more suitable for the Yelp dataset.
-Table 4. Detailed settings of different $\bm{\lambda}$.
-Tag Type Function
-$\lambda_{\text{f1}}$ $\lambda_{\text{fix}}$ 0.2
-$\lambda_{\text{f2}}$ $\lambda_{\text{fix}}$ 0.1
-$\lambda_{\text{f3}}$ $\lambda_{\text{fix}}$ 0.003
-$\lambda_{\text{d1}}$ $\lambda_{\text{decline}}$$\begin{cases}0.2-0.2*\text{ep}/20+3\times 20^{-3}&0<\text{ep}<20\\ 3\times 10^{-3}&\text{ep}>=20\\ \end{cases}$
-$\lambda_{\text{d2}}$ $\lambda_{\text{decline}}$$\begin{cases}0.2-0.2*\text{ep}/30+3\times 10^{-3}&0<\text{ep}<30\\ 3\times 10^{-3}&\text{ep}>=30\\ \end{cases}$
-$\lambda_{\text{d3}}$ $\lambda_{\text{decline}}$$\begin{cases}0.2-0.2*\text{ep}/40+3\times 10^{-3}&0<\text{ep}<40\\ 3\times 10^{-3}&\text{ep}>=40\\ \end{cases}$
-$\lambda_{\text{d4}}$ $\lambda_{\text{decline}}$$\begin{cases}0.2-0.2*\text{ep}/80+3\times 10^{-3}&0<\text{ep}<80\\ 3\times 10^{-3}&\text{ep}>=80\\ \end{cases}$
-$\lambda_{\text{s1}}$ $\lambda_{\text{smooth}}$$\begin{cases}2*(1-\text{ep}/150)^{6}+3\times 10^{-3}&0<\text{ep}<150\\ 3\times 10^{-3}&\text{ep}>=150\\ \end{cases}$
-$\lambda_{\text{s2}}$ $\lambda_{\text{smooth}}$$\begin{cases}2*(1-\text{ep}/150)^{9}+3\times 10^{-3}&0<\text{ep}<150\\ 3\times 10^{-3}&\text{ep}>=150\\ \end{cases}$
-$\lambda_{\text{s3}}$ $\lambda_{\text{smooth}}$$\begin{cases}0.05*(1-\text{ep}/150)^{10}+3\times 10^{-3}&0<\text{ep}<150\\ 3\times 10^{-3}&\text{ep}>=150\\ \end{cases}$
-Figure 10. Illustrations of different $\bm{\lambda}$.
-Table 5. The comparison performance of different $\bm{\lambda}$ settings. Bold scores represent the highest results.
-Dataset Metric $\lambda_{\text{fix}}$ $\lambda_{\text{decline}}$ $\lambda_{\text{smooth}}$
-$\lambda_{f1}$ $\lambda_{f2}$ $\lambda_{f3}$ $\lambda_{d1}$ $\lambda_{d2}$ $\lambda_{d3}$ $\lambda_{d4}$ $\lambda_{s1}$ $\lambda_{s2}$ $\lambda_{s3}$
-Beauty HR@5 0.0453 0.0507 0.0493 0.0320 0.0426 0.0479 0.0584 0.0617 0.0630 0.0564
-HR@10 0.0643 0.0704 0.0669 0.0471 0.0636 0.0688 0.0844 0.0855 0.0859 0.0795
-NDCG@5 0.0318 0.0366 0.0357 0.0224 0.0298 0.0340 0.0414 0.0435 0.0449 0.0399
-NDCG@10 0.0379 0.0429 0.0413 0.0272 0.0365 0.4065 0.0498 0.0514 0.0522 0.0473
-Toys HR@5 0.0470 0.0582 0.0458 0.0113 0.0284 0.0396 0.0676 0.0684 0.0690 0.0594
-HR@10 0.0627 0.0787 0.0648 0.0195 0.0418 0.0552 0.0889 0.0901 0.0903 0.0804
-NDCG@5 0.0352 0.0436 0.0329 0.0083 0.0203 0.0287 0.0503 0.0510 0.0518 0.0447
-NDCG@10 0.0402 0.0502 0.0390 0.0109 0.0244 0.0337 0.0571 0.0579 0.0587 0.0514
-Yelp HR@5 0.0241 0.0269 0.0247 0.0344 0.0405 0.0396 0.0251 0.0327 0.0289 0.0398
-HR@10 0.0363 0.0411 0.0424 0.0547 0.0611 0.0605 0.0174 0.0514 0.0192 0.0591
-NDCG@5 0.0169 0.0179 0.0163 0.0232 0.0272 0.0266 0.0388 0.0223 0.0443 0.0270
-NDCG@10 0.0207 0.0225 0.0220 0.0297 0.0338 0.0333 0.0218 0.0282 0.0241 0.0333
-E.2. Other Hyper-parameters Analysis
-Figure 11. Hyperparameters analysis of our DCRec  on Beauty dataset.
-We report the experimental results for key hyperparameters on the Beauty dataset, including the noise schedule $\beta$, number of diffusion steps $T$, and scaling factor $\delta$.
-In detail, $\beta$ is tuned within $\{0.01,0.02,0.04,0.06,0.08,0.1\}$, diffusion steps $T$ is tuned within $\{5,10,15,20,30,50,75,100,150,200\}$ and scaling factor $\delta$ is tuned within $\{0.1,0.2,0.5,0.75,1\}$.
-As shown in Figure 11, we can find for the noise schedule $\beta$, performance remains stable for values up to 0.04 but sharply declines at 0.06, indicating that excessive noise negatively impacts the model.
-The number of training diffusion steps $T$ improves performance up to 20 steps, after which it plateaus, suggesting that 20 steps are sufficient.
-The scaling factor $\delta$ performs best in the range of 0.2 to 0.5, with both smaller and larger values resulting in decreased performance.
-Appendix F Implementation Details
-We implement our model and baselines using Pytorch (Paszke et al., 2019) and conduct our experiments on an NVIDIA RTX 3090 GPU with 24G memory.
-For a fair comparison, we carefully tuned the model parameters of all the baselines on the validation set.
-Then we choose the model parameters that achieve the best performance on the validation set to compare.
-For our DCRec, based on the hyperparameter analysis, we set the batch size to 512, the item embedding dimension to 128, learning rate of $34\times 1e-$ 3, maximum sequence length to 50, $\beta$ to 0.02, the dropout rate to 0.1, the embedding dropout to 0.4, the hidden factor to 128, $\delta$ to 0.1, the number of transformer blocks to 4, and the number of diffusion steps $T$ to 50.
-For $\lambda$, we adopt $\lambda_{smooth}=\lambda_{s2}$ for the Beauty and Toys datasets, while for the Yelp dataset, we use $\lambda_{decline}=\lambda_{d2}$, as determined by the results in Table 5 result.
-We choose Adam (Kingma and Ba, 2015) as our optimization strategy and use early stopping with a patience of 5 to prevent overfitting.
+Therefore, the strategy of partially noisy input is unfeasible for DCRec.
+Figure 9. The impact of history conditioning with noise and without noise on Beauty and Toys datasets, reflecting by HR@5 and NDCG@5.
+5. Related Work
+5.1. Diffusion Models
+Diffusion models, derived from non-equilibrium thermodynamics, have achieved notable success in areas like computer vision, sequence modeling, and audio processing (Yang et al., 2024b).
+Two primary paradigms exist: unconditional models and conditional models.
+Unconditional models, such as DDPM (Ho et al., 2020), aim to push the performance boundaries of generative models.
+Conditional models incorporate guidance from labels to control generated samples.
+SGGM (Yang et al., 2022) conditions on self-produced labels, while You et al.
+(You et al., 2023) demonstrate how few labels and pseudo-training enhance both large-scale diffusion models and semi-supervised learners.
+Dhariwal (Dhariwal and Nichol, 2021) uses classifier guidance to improve sample quality, and Ho and Salimans (Ho and Salimans, 2022b) combine conditional and unconditional scores to balance sample quality and diversity.
+Multi-modal guidance further enriches the diffusion process, as seen in DiffuSeq (Gong et al., 2023) for NLP tasks and SDEdit (Meng et al., 2022) for image translation, with latent diffusion models (LDM) (Rombach et al., 2022) offering flexible, unified semantics.
+unCLIP (Ramesh et al., 2022) and ConPreDiff (Yang et al., 2023a) integrate CLIP latents in text-to-image generation.
+5.2. Diffusion for Sequential Recommendation
+Diffusion models show promise in the sequential recommendation by using user behavior sequences as guidance in the denoising process (Xie et al., 2024; Yang et al., 2023b; Li et al., 2024b).
+DiffuRec (Li et al., 2024b) corrupts target item embeddings with Gaussian noise and uses a Transformer to iteratively reconstruct them.
+DreamRec (Yang et al., 2023b) generates the oracle next-item embeddings based on user preferences but struggles with scalability due to its lack of negative sampling.
+While promising to some extent, these implicit conditional methods may oversimplify complex user behaviors by compressing them into a single vector, potentially losing valuable sequential patterns and contextual information.
+Additionally, DCDR (Lin et al., 2024a) adopts a discrete diffusion framework for progressive re-ranking, generating a sequence rather than a specific target item.
+6. Conclusion
+In this paper, we proposed the DCRec  to address key limitations in existing diffusion-based sequential recommender systems.
+We first propose the Dual Conditional Diffusion Models for the sequential recommendation, introducing a novel diffusion framework for the SR.
+Based on this framework, we propose the DCDT  which can generate accurate target item embedding based on both implicit and explicit conditions.
+Besides, our framework simplifies the diffusion process by reducing some inference steps with high similarity.
+Extensive experiments on public datasets validate the effectiveness of DCRec, demonstrating its superiority over state-of-the-art methods in both recommendation accuracy and computational efficiency.